### v1

# Iterative Utility Judgment Framework via LLMs Inspired by Relevance in Philosophy

Abstract
Utility and topical relevance are critical measures in information retrieval (IR), reflecting system and user perspectives, respectively. While topical relevance has long been emphasized, utility is a higher standard of relevance and is more useful for facilitating downstream tasks, e.g., in Retrieval-Augmented Generation (RAG). When we incorporate utility judgments into RAG, we realize that the topical relevance, utility, and answering in RAG are closely related to the three types of relevance that Schutz discussed from a philosophical perspective. They are topical relevance, interpretational relevance, and motivational relevance, respectively. Inspired by the dynamic iterations of the three types of relevance, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to promote each step of the cycle of RAG. We conducted extensive experiments on multi-grade passage retrieval and factoid question-answering datasets (i.e., TREC DL, WebAP, and NQ). Experimental results demonstrate significant improvements in utility judgments, ranking of topical relevance, and answer generation upon representative baselines, including multiple single-shot utility judging approaches. Our code and benchmark can be found at https://anonymous.4open.science/r/ITEM-B486/.
Iterative Utility Judgment Framework via LLMs Inspired by Relevance in Philosophy
Hengran Zhang1,2 Keping Bi1,2 Jiafeng Guo1,2 Xueqi Cheng1,2 1CAS Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences; 2School of Computer Science and Technology, University of Chinese Academy of Sciences; {zhanghengran22z, bikeping, guojiafeng, cxq}@ict.ac.cn,
1 Introduction
Utility and topical relevance measure Information Retrieval (IR) performance from the views of users and systems, respectively Bruce (1994). Utility refers to the usefulness of retrieval items to an information seeker, gauged by their “value”, “appropriateness in resolution of a problem” or “reduction of uncertainty”, etc. Saracevic (1996); Schamber and Eisenberg (1988); Saracevic (1975). In contrast, topical relevance measures the matching between the subjects of queries and documents, which indicates aboutness Saracevic et al. (1988). Topical relevance does not necessarily mean utility, while utility indicates a high standard of relevance.
Since topical relevance is relatively easy to observe and measure Schamber et al. (1990), the studies of IR models have been primarily focused on improving the system view of relevance for a long time Bruce (1994). In recent years, most IR models have been experimented on the MSMARCO dataset Nguyen et al. (2016) and optimized towards the binary labels, which is inadequate to capture utility. More recently, leveraging Large Language Models (LLMs) OpenAI (2022); Meta (2024); Jiang et al. (2023a); OpenAI (2023) for topical relevance judgments Khramtsova et al. (2024); Faggioli et al. (2023); Thomas et al. (2024); Sun et al. (2023) has drawn more attention than judging utility Zhang et al. (2024).
In the modern LLM era, Retrieval-Augmented Generation (RAG) has become a hot research topic that facilitates LLMs with external knowledge Xie et al. (2023); Shi et al. (2023); Izacard et al. (2023); Su et al. (2024); Glass et al. (2022). It requests retrieval results useful for answering the question. Given the constrained bandwidth of LLM inputs, it is essential to prioritize high-value results to guide LLMs. Consequently, utility needs to be emphasized more than topical relevance in RAG. In this paper, we aim to promote the utility judgment performance of LLMs so that relevance ranking can be boosted and RAG can be enhanced by high-utility references.
We find that the interpretation of relevance in philosophy by Schutz Schutz (1970) is closely related to RAG. Schutz considered relevance as the property that determines the connections and relations in our lifeworld. He identified three types of basic and interdependent relevance that interact dynamically within a “system of relevancies” Saracevic (1996): (i) Topical relevance, which refers to the perception of what is separated from one’s experience to form one’s present object of concentration; (ii) Interpretational relevance, which involves the past experiences in understanding the currently concerned object; and (iii) Motivational relevance, which refers to the course of action to be adapted based on the interpretations. The motivational relevance, in turn, helps obtain additional materials to become a user’s new experience, which further facilitates topical and interpretational relevance. Schutz posited that one’s perception of the world may be enhanced by this dynamic interaction, as shown in Figure 1. We consider that topical relevance, interpretational relevance, and motivational relevance are closely related to topical relevance, utility, and answering respectively in RAG. Inspired by the dynamic interaction of the relevance system, we believe the dynamic interactions between the three components in RAG can promote the performance of each step.
To verify the idea, we leverage LLMs to perform each step in RAG shown in Figure 1, i.e., the judgments of topical relevance, utility, and answer generation. While previous work Zhang et al. (2024) has shown that utility judgments via LLMs can be enhanced by simultaneously answering the question, iteratively conducting utility judgments and question answering (QA) has not been explored, not to mention that topic relevance can also be involved in the iteration. We propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps.
We experiment on various information-seeking tasks, i.e., multi-grade passage retrieval on TREC DL Craswell et al. (2020), multi-grade non-factoid answer passage retrieval on WebAP Yang et al. (2016), and factoid QA Kwiatkowski et al. (2019). For multi-grade passage retrieval, we consider the ones with the highest grade to be of utility and we focus on the performance of utility judgments and topical relevance ranking. For factoid QA, we emphasize the answer accuracy. Experimental results have demonstrated that ITEM can significantly outperform competitive baselines including various single-shot judgment approaches in terms of utility judgments, ranking of topical relevance, and answer generation, which confirms the viability of adaptation of Schutz’s viewpoint of the relevance system into RAG.
2 Related Work
Relevance-Focused Information Retrieval. Relevance is a crucial concept in information retrieval (IR), and the automatic generation of relevance judgments has been a long-standing goal in IR that has been studied over the past decades Saracevic (1975); Järvelin and Kekäläinen (2017); Makary et al. (2017, 2016); MacAvaney and Soldaini (2023). Recent studies demonstrate promising outcomes in utilizing LLMs to automatically generate relevance judgments Khramtsova et al. (2024); Faggioli et al. (2023); Thomas et al. (2024) and mainly contain the following three approaches, i.e., pointwise Zhuang et al. (2024); Nogueira and Cho (2019), pairwise Qin et al. (2023); Jiang et al. (2023b), and listwise Sun et al. (2023); Pradeep et al. (2023). For example, Faggioli et al. (2023) used zero/few-shot prompting to instruct LLMs to judge the relevance of an item to a query in a pointwise approach. Sun et al. (2023) analyzed the capabilities of LLMs in the relevance ranking.
Utility-Focused Information Retrieval. Utility is a distinct measure of relevance compared to topical relevance Zhao et al. (2024); Saracevic et al. (1988); Saracevic (1975, 1996); Ji et al. (2024); Zhang et al. (2023), and more recently, Zhang et al. (2024) highlighted the use of LLMs for utility judgments. However, Zhang et al. (2024) only conducted a preliminary exploration of LLMs in utility judgments. Our work aims to further explore how to improve the performance of utility judgments for LLMs.
Retrieval-Augmented Generation (RAG). RAG approaches are widely employed to mitigate the hallucination issues in large language models (LLMs) Xie et al. (2023); Zhou et al. (2024); Su et al. (2024). The current RAG approaches are categorized as follows: (i) single-round retrieval Borgeaud et al. (2022); Lewis et al. (2020); Glass et al. (2022); Izacard et al. (2023); Shi et al. (2023), which involves using the initial input as a query to retrieve information from an external corpus and then the information is incorporated as part of the input for the model; and (ii) multi-round retrieval Su et al. (2024); Jiang et al. (2023c); Ram et al. (2023); Khandelwal et al. (2020); Trivedi et al. (2023), which need multi-round retrieval based on feedback from LLMs.
Iterative Relevance Feedback via LLMs. Recent works Li et al. (2023); Shao et al. (2023) have achieved great success in using LLMs to obtain the information needs of the question as pseudo-relevance feedback for iterative retrieval. They posit that a single retrieval may not yield comprehensive information, thus requiring multiple retrievals. In contrast, our methodology involves making iterative utility judgments on the results obtained from a single retrieval. Given the substantial operational costs associated with retrieval systems, the expense incurred from conducting multiple retrievals for a single query becomes even more prohibitive.
3 Utility Judgments (UJ) via LLMs
We formally define the task of utility judgments via LLMs and propose an Iterative utiliTy judgmEnt fraMework (ITEM) for utility judgments.
3.1 Notations and Definitions
Given a question and a list of retrieved passages . The goal of utility judgments for LLMs is to identify a set of passages that have utility in answering the question . There are two typical input approaches for LLMs:
Listwise, , where is the instruction of LLMs and is the utility judgments function based on the whole candidate list and the output is a set of passages.
Pointwise, where is the instruction of LLMs and is the utility judgments function with a single passage as the input.
3.2 Sing-Shot Utility Judgments
The most common approach to judge utility is to perform a single-shot utility judgment, i.e., , where is the utility judgments function for LLMs, which can be or . Zhang et al. (2024) proposed to generate a pseudo-answer while conducting utility judgments, which can help LLMs to judge utility better, i.e., .
3.3 Iterative utiliTy judgmEnt fraMework (ITEM)
Schutz Schutz (1970) emphasized the existence of various types of relevance and underscored the interactivity and interdependence between these various types from a much broader arena than information science. Inspired by the powerful insight, we propose an Iterative utiliTy judgmEnt fraMework (ITEM). Specifically, as shown in Figure 4, dynamic iteration of the topical relevance, pseudo answer generation, and utility judgments.
ITEM with Answering in the Loop (ITEM-A).
In a formal sense, at iteration (), the pseudo answer is generated based on the judging result in the previous iteration and then judged based on the and candidate passages list to get a set of passages that have utility :
where is an answer generation function, is a utility judgment function, which can be or , is the answer prompts for LLMs (as detailed in Figure 2), and is the utility judgments prompts for LLMs (as detailed in Figure 3). can be in two forms: (i) explicit answerto the question ; (ii) implicit answerthat specifies necessary information to answer the question . We consider as the initial candidate set and is the initial results ranked by a retriever like BM25 Robertson et al. (2009).
ITEM with both Answering and Ranking of Topical Relevance in the Loop (ITEM-AR). In Schutz’s theory, topical relevance, motivational relevance, and interpretive relevance are all dynamically affected. In the ITEM-A framework, the topical relevance is not updated during the iteration process. Consequently, we have incorporated a relevance ranking task into the ITEM framework, which ensures that all three tasks are executed in a loop.
Formally, at iteration (), the answer is generated based on the judging result in the previous iteration. Subsequently, given , the passage list is ranked based on the relevance to the question to obtain . Finally, the judging result is derived using the input and the answer :
where and are the relevance ranking function and the relevance ranking prompt for LLMs (as detailed in Figure 5), respectively.
At iteration , we have two ways to produce the set : (i) asking LLMs to identify the set of passages that have utility, which we call ITEM-As or ITEM-ARs; (ii) asking LLMs to provide a ranked passage list with the listwise input approach and we deem the set of first elements in the list as , which we call ITEM-Ar or ITEM-ARr. We set as 5 and more details of selection are shown in Appendix B.4. In our experiments, we find that ITEM-ARr does not improve ranking performance as well as ITEM-Ar (see Section 6.1 for experimental analysis), so we do not use ITEM-ARr in our experiments to obtain .
Full details of all prompts can be found in Appendix A. We stop the iteration when at most iterations are reached or the set of passages considered to have utility does not change, i.e., or . If the value of is relatively small, it may negatively impact the performance of the iteration. Conversely, if the value of is too large, it may increase the computational cost of the iteration. Considering the balance between performance and efficiency, we selected =3. Further discussion about is provided in Section 6.1.
4 Experimental Setup
4.1 Datasets
Our experiments are conducted on three benchmark datasets, including, two retrieval datasets, i.e., TREC DL Craswell et al. (2020), WebAP Yang et al. (2016), and an open-domain question answer (ODQA) dataset, i.e., NQ Kwiatkowski et al. (2019).
TREC. We use the TREC-DL19 and TREC-DL20 datasets Craswell et al. (2020). Judgments of TREC DL are on a four-point scale, i.e., “perfectly relevant”, “highly relevant”, “related”, and “irrelevant”. We consider the passages that are “perfectly relevant” to have utility. We filter questions of two datasets that contain the passages labeled by the highest point and combine them to form a whole dataset, i.e., the TREC dataset in our experiments.
WebAP. WebAP Yang et al. (2016) is a non-factoid answer passage collection built on Gov2. More details about non-factoid answer passage retrieval are provided in Appendix F. It retrieves the top 50 documents with the Sequential Dependency Model (SDM) Metzler and Croft (2005) for the description queries that need to be answered with passages. Relevant documents are annotated, resulting in 3843 passages from 1200 documents, categorized as “perfect”, “excel”, “good”, and “fair”. In our experiments, we split the remaining documents into passages of 2 or 3 sentences to be irrelevant and considered “perfect” to have utility.
NQ. Natural Questions(NQ) consist of factoid questions issued to the Google search engine Kwiatkowski et al. (2019). Each question is annotated with a long answer (typically a paragraph) and a short answer (one or more entities). Following Zhang et al. (2024), we use the questions that have long answers in our experiments. Detailed statistics of all the experimental datasets are shown in Table 16 in the Appendix E.
4.2 Evaluation metrics
For the utility judgments task, we evaluate the results of judgments using Precision, Recall, and F1. For the ranking task, we use the normalized discounted cumulative gain (NDCG) Järvelin and Kekäläinen (2017) to evaluate the ranking performance. For the answer generation task, we use the standard exact match metric (EM) and F1, as used by Rajpurkar et al. (2016).
4.3 LLMs
we conduct our experiments using several representative LLMs, i.e., (i) the well-known ChatGPT OpenAI (2022) (we use the gpt-3.5-turbo-1106 version), (ii) Mistral Jiang et al. (2023a) (the Mistral-7B-Instruct-v0.2 version after instruction tuning), and (iii) Llama 3 Meta (2024) (the Meta-Llama-3-8B-Instruct version promoted with instruction tuning based on the vanilla Llama 3). To ensure the reproducibility of the experiments, the temperature for all experiments is set to 0.
4.4 Retrievers
We use two representative retrievers to gather candidate passages in for utility judgments. Following with previous works Zhang et al. (2024); Sun et al. (2023), we employ RocketQAv2 Ren et al. (2021) and BM25 Robertson et al. (2009) for the NQ dataset and retrieval datasets(i.e., TREC and WebAP datasets), respectively. Based on the retrieval results to build the we have two settings: (i) For TREC and WebAP datasets, we select the top-20 BM25 retrieval results. If these do not include passages with utility, we replaced the last one with a utility-annotated passage. (ii) For the NQ dataset, we use the top-10 dense retrieval results to form the candidate list , following the GTU setting of Zhang et al. (2024).
4.5 Baselines
We choose the following baselines for comparison:
Sing-shot utility judgments: (i) Vanilla: Ask LLMs to provide utility judgments based on the instruction directly. (ii) UJ-ExpA: Utility judgments facilitated with explicit answers, which is shown to be effective in Zhang et al. (2024). (iii) UJ-ImpA: We utilize the necessary information that is necessary for answering the question to help with the utility judgments task. More details can be found in Section 3.2.
-sampling is proposed by Zhang et al. (2024): In addition to the original input order, -sampling method randomizes the order of input passage list times and aggregates the utility judgement results through voting. The output of the -sampling method each time contains explicit answers and utility judgments. Our iterative framework outputs both answers and utility judgments at each iteration. To ensure fairness in calculating the parameters of LLM, we set . The detailed comparison of the parameter calculations of the two is in Appendix G. The -sampling method requires outputs of LLMs, which can be more expensive for practical applications involving the LLMs. Moreover, it is difficult to determine the stability of the utility judgments due to the sampling technique.
| Method | TREC | WebAP | ||||||||||
| Listwise | Pointwise | Listwise | Pointwise | |||||||||
| Mis. | Lla. | Cha. | Mis. | Lla. | Cha. | Mis. | Lla. | Cha. | Mis. | Lla. | Cha. | |
| Vanilla | 45.67 | 49.39 | 55.19 | 45.11 | 47.64 | 49.84 | 20.79 | 21.79 | 28.43 | 23.05 | 25.09 | 26.85 |
| UJ-ExpA | 54.10 | 52.83 | 57.49 | 43.53 | 53.73 | 48.09 | 27.94 | 26.99 | 30.50 | 25.27 | 29.25 | 27.44 |
| UJ-ImpA | 48.29 | 48.22 | 56.18 | 48.31 | 50.20 | 48.83 | 25.06 | 26.22 | 29.89 | 28.35 | 25.29 | 26.32 |
| -sampling | 52.31 | 52.68 | 60.49 | - | - | - | 30.16 | 28.97 | 31.49 | - | - | - |
| ITEM-As ExpA (1) | 53.78 | 53.66 | 62.52 | 49.44 | 52.09 | 53.61 | 29.76 | 27.50 | 36.89 | 29.10 | 31.08 | 32.02 |
| ITEM-As ExpA (3) | 54.86 | 56.03 | 63.18 | 51.74 | 52.46 | 55.74 | 31.65 | 29.32 | 39.57 | 30.50 | 32.67 | 31.43 |
| ITEM-As ImpA (1) | 49.39 | 53.73 | 58.11 | 46.01 | 53.68 | 54.61 | 26.06 | 25.59 | 34.97 | 28.28 | 30.53 | 29.34 |
| ITEM-As ImpA (3) | 52.05 | 55.14 | 60.56 | 46.59 | 53.76 | 54.90 | 28.36 | 26.10 | 40.78 | 30.13 | 29.64 | 32.54 |
| ITEM-ARs (1) | 52.34 | 48.97 | 62.00 | - | - | - | 35.50 | 31.44 | 36.58 | - | - | - |
| ITEM-ARs (3) | 56.27 | 52.10 | 61.37 | - | - | - | 37.06 | 29.08 | 38.58 | - | - | - |
5 Experimental Results of LLMs
Our framework consists of three tasks: utility judgments, ranking of topical relevance, and answer generation. This section will present the performance of each task within our framework. By default, the pseudo answer is the explicit answer in all experiments, if not specified otherwise.
5.1 Utility Judgments on Retrieval Datasets
Tabel 1 shows the F1 performance on the TREC and WebAP datasets using three LLMs. Since ITEM-Ar has poor F1 performance in utility judgments (refer to Table 12 in Appendix B.1 for details), we restrict our experiments to ITEM-As and ITEM-ARs in this section.
| Method | Mistral | Llama 3 | ChatGPT | |||
| T | W | T | W | T | W | |
| BM25 | 58.69 | 21.89 | 58.69 | 21.89 | 58.69 | 21.89 |
| RankGPT | 69.81 | 29.34 | 75.61 | 41.73 | 80.56 | 42.49 |
| ITEM-ARs (1) | 71.29 | 37.48 | 77.22 | 43.80 | 81.38 | 48.42 |
| ITEM-ARs (3) | 73.24 | 45.45 | 74.80 | 44.87 | 82.89 | 48.80 |
| ITEM-Ar (1) | 70.57 | 37.11 | 73.95 | 40.89 | 80.79 | 50.30 |
| ITEM-Ar (3) | 74.27 | 43.80 | 77.34 | 45.88 | 83.12 | 51.61 |
ITEM with a Single Iteration vs. Baselines. All three LLMs using our ITEM with a single iteration generally outperform the single-shot utility judgments on both datasets and may even surpass the -sampling method. For example, ChatGPT on the TREC dataset using our ITEM-As ExpA and ImpA in the listwise approach improve the F1 performance by 8.7% and 3.4% over UJ-ExpA and UJ-ImpA, respectively. On the WebAP dataset, Mistral and Llama 3 use single-iteration ITEM-ARs to outperform all baselines and ITEM-As. This might be due to that better topical relevance ranking leads to better utility judgments, which is also better than -sampling in terms of efficacy and efficiency.
ITEM with Multiple Iterations vs. ITEM with Single Iteration. All LLMs using our ITEM-A and ITEM-AR generally demonstrate improved performance with multiple iterations compared to single iterations. For instance, on the WebAP dataset, Mistral, Llama 3, and ChatGPT (using our ITEM-A ExpA) improved their F1 scores in the listwise approach by 6.4%, 6.6%, and 7.3%, respectively, after multiple iterations. However, in some specific cases, multiple iterations may not outperform single iterations, likely due to the unpredictable nature of zero-shot settings. Moreover, our method achieves state-of-the-art performance compared to all baselines by leveraging the iterative framework. The performance improvement from multiple iterations underscores the significance of iterative interaction and further supports Schutz’s interactive framework.
ExpA vs. ImpA. For the single-shot baselines, explicit answers generally outperform implicit answers in utility judgments when using the listwise approach, while the opposite is true for the pointwise approach. This may be because pointwise inputs provide less information, making explicit answers less effective, whereas implicit answers better address the question’s information needs. Unlike the pseudo answer generated by the input passage of the single-shot pointwise approach, our ITEM’s pointwise approach generates the pseudo answer based on the results of utility judgments from the previous round. Therefore, after multiple iterations, ITEM shows greater improvement with explicit answers compared to implicit answers in both input approaches in most cases.
Comparisons between LLMs. ChatGPT outperforms other LLMs on both datasets using both input approaches. For instance, on the TREC dataset, ChatGPT achieves an F1 improvement of 15.2% and 12.8% using the listwise approach after multiple iterations of ITEM-As ExpA method, compared to Mistral and Llama 3, respectively.
Listwise vs. Pointwise. The average performance of utility judgments for LLMs is better with the listwise approach than with the pointwise approach. For example, Mistral shows a 6.0% F1 improvement with the listwise approach over the pointwise approach after multiple iterations of the ITEM-As ExpA method on the TREC dataset. More detailed performance metrics, i.e., Precision, Recall, F1, are provided in Appendix B.1.
5.2 Ranking Performance on Retrieval Datasets
We also assess whether the ranking of topical relevance has been improved within ITEM. In terms of topical relevance, the utility ranking results produced by ITEM-Ar and the relevance ranking results yielded from ITEM-ARs. From Table 2, we can observe that: (i) Our ITEM with single iteration significantly improves the ranking of topical relevance performance compared to the RankGPT. For instance, relevance ranking outperforms RankGPT in NDCG@5 by 2.1% on the TREC dataset and 27.7% on the WebAP dataset using the Mistral. The performance improvement may stem from the interaction between tasks. (ii) After iterations, relevance ranking, and utility ranking performances have been improved on both datasets and all LLMs. The topical relevance ranking benefits from our dynamic iterative framework, confirming Schutz’s theory of dynamic iterative interaction. (iii) Experiments show that utility ranking generally outperforms relevance ranking, which confirms the effectiveness of utility in the ranking of topical relevance.
5.3 Results of Answer Generation
From Table 1, we find that the listwise approach generally outperforms the pointwise approach for utility judgments. Consequently, our NQ dataset experiments utilize only the listwise approach.
| References | Mistral | Llama 3 | ChatGPT | |||
| EM | F1 | EM | F1 | EM | F1 | |
| Golden | 46.09 | 62.59 | 64.45 | 76.64 | 66.40 | 76.86 |
| RocketQAv2 | 31.58 | 47.69 | 50.96 | 62.01 | 46.54 | 57.00 |
| Vanilla | 31.16 | 47.43 | 49.09 | 60.56 | 48.52 | 58.64 |
| UJ-ExpA | 32.76 | 48.46 | 49.63 | 61.10 | 47.72 | 58.01 |
| UJ-ImpA | 30.67 | 46.83 | 48.88 | 60.26 | 49.01 | 59.30 |
| 5-sampling | 33.24 | 48.84 | 48.72 | 60.71 | 48.90 | 58.97 |
| ITEM-As (1) | 32.98 | 49.00 | 50.16 | 61.88 | 49.38 | 59.78 |
| ITEM-As (3) | 33.73 | 49.63 | 50.27 | 62.09 | 49.69 | 60.18 |
| ITEM-ARs (1) | 33.30 | 49.26 | 50.27 | 61.69 | 49.52 | 59.64 |
| ITEM-ARs (3) | 33.40 | 49.27 | 49.36 | 60.97 | 49.06 | 59.67 |
The following observations can be made from Table 3: (i) The -sampling method has the best performing, especial on F1 score, for all the baselines. (ii) ITEM outperforms baselines on all metrics on all LLMs (except for the EM score of Llama 3), indicating that ITEM can help the LLMs to find better evidence for generating answers. (iii) The results from Llama 3 have better answer performance than ChatGPT, and we suspect that Llama 3 has been over-tuned on the NQ dataset.
6 Further Analyses
6.1 Effect of Iteration Numbers
Figure 6 illustrates the performance of the LLMs in making utility judgments under ITEM-As with varying maximum iteration rounds on the TREC dataset. Figure 7 illustrates the ranking of topical relevance performance in terms of values. Since the NQ dataset lacks annotations for passage utility, we automatically annotated the retrieved passages based on the utility standard, as detailed in Appendix D. Table 4 displays the results for utility judgments and answer generation using Mistral under ITEM-As on the NQ dataset.
| ITEM-As | ITEM-ARs | |||||
| Utility | Answer | Utility | Answer | |||
| F1 | EM | F1 | F1 | EM | F1 | |
| =1 | 28.83 | 32.98 | 49.00 | 24.96 | 33.30 | 49.26 |
| =2 | 29.02 | 34.31 | 50.08 | 25.87 | 33.57 | 49.16 |
| =3 | 28.90 | 33.73 | 49.63 | 26.17 | 33.40 | 49.27 |
| =4 | 29.01 | 34.21 | 50.07 | 26.54 | 33.46 | 49.24 |
| =5 | 28.91 | 33.78 | 49.63 | 26.48 | 33.89 | 49.58 |
As can be seen from Figure 6, Figure 7 and Table 4, varying the value of affects the performance of utility judgments, ranking of topical relevance and answer generation. However, the performance when is consistently better than when . Moreover, utility ranking generally outperforms relevance ranking, which confirms the effectiveness of utility in the ranking of topical relevance. Further details on the experiments with different values are provided in Appendix B.1.
In addition, we evaluate two ranking performances of ITEM-ARr, with the experimental results shown in Table 5. Under the ITEM-ARr framework, relevance ranking is followed by utility ranking, resulting in further improved utility ranking performance compared to relevance ranking. However, as seen in Table 2 and Table 5, performing ranking twice in the same iteration may not yield better ranking results than performing utility ranking once in the iteration.
| NDCG@5 | NDCG@10 | NDCG@20 | Utility-F1 | |
| 1 | 71.29 / 72.77 | 72.90 / 74.96 | 84.56 / 85.75 | 43.13 |
| 2 | 72.54 / 70.99 | 74.81 / 73.76 | 85.77 / 85.28 | 40.21 |
| 3 | 72.07 / 74.14 | 74.14 / 76.63 | 85.53 / 86.57 | 45.67 |
| 4 | 71.02 / 71.06 | 74.30 / 74.03 | 85.09 / 85.16 | 43.82 |
| 5 | 72.26 / 70.12 | 75.83 / 72.59 | 85.88 / 84.77 | 44.10 |
6.2 Iteration Stop Conditions
The stopping condition takes into account the answer performance in our ITEM in addition to the utility judgments. we compute the ROUGE-L Lin (2004) score of the answer in two iterations, and stop the iteration if the ROUGE-L of and is greater than . The utility judgments performance of different iteration stop conditions are shown in Figure 8. More details of experiments on iteration stop conditions are shown in Appendix B.2.
The results show that using different stopping conditions does affect the performance of utility judgments. However, using the answer as a stopping condition, different models on different datasets may need to look for different , which is not very flexible.
7 Case Study
Table 16 and Table 17 in Appendix C present two good cases (demonstrating improved utility judgments after multiple iterations) from the TREC dataset using Mistral on ITEM-As. In Table 16, the first pseudo answer, though relatively correct, includes irrelevant information, leading to a misjudgment of “Passage-2” as “utility”. “Passage-2” does not address the question of what tissue bronchioles belong to but instead discusses the type of muscles that control the size of the bronchioles in the lungs. Based on the results of the first round of utility judgments, the second round of the pseudo answer is more accurate and free of irrelevant content. Consequently, all three passages in the second round of utility judgments have utility in answering the question. In Table 17, the first pseudo answer is correct, but two irrelevant passages are judged as “utility”. The second pseudo answer, with slight rewording, results in all passages being correctly identified as “utility”.
8 Conclusion
In this paper, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps, inspired by Schutz’s philosophical discussion of relevance. Our framework achieves state-of-the-art performance in zero-shot scenarios, outperforming previous methods in utility judgments, ranking of topical relevance, and answer generation tasks. Our experiments also highlight the significance of dynamic interaction in achieving high performance and stability. Future directions include developing better fine-tuning strategies for utility judgments and creating end-to-end solutions for retrieval and utility judgments.
Limitations
In this paper, we proposed an iterative utility judgment framework inspired by relevance in philosophy. There are three primary limitations that should be acknowledged: (i) Our methods are applied in zero-shot scenarios without any training. The zero-shot method has inherent disadvantages, e.g., being sensitive to prompts and having unstable performance. The zero-shot approach itself does not enhance the LLMs’s inherent capability in utility judgments but rather employs strategies to improve performance on utility judgment tasks. Future research should explore designing more effective training methods, e.g., utilizing our iterative framework with self-evolution techniques Singh et al. (2023), to genuinely enhance the LLMs’s ability in utility judgments through training. (ii) The number of candidate passages in the search scenario is much larger than 20. The number of search results we assumed is too small. We need to continue to study utility judgments in large-scale scenarios in the future. (iii) Although our iterative framework is very effective, the number of iterations will increase the cost of calling large models. How to reduce the iteration cost is also something that needs further improvement in the future.
Ethics Statement
In conducting this research, we have prioritized ethical considerations at every stage to ensure the responsible development and application of AI technologies. Our research does not rely on personally identifiable information or require manually annotated datasets. We firmly believe in the principles of open research and the scientific value of reproducibility. To this end, we have made all models, data, and code associated with our paper publicly available on GitHub. This transparency not only facilitates the verification of our findings by the community but also encourages the application of our methods in other contexts.
Appendix A Instruction Details
A.1 Instruction of Listwise and Pointwise Approaches
For the prompts of the NQ dataset using ChatGPT, we follow the setting of Zhang et al. (2024), otherwise, we use the following prompts. Following Sun et al. (2023), we input N passages using the form of multiple rounds of dialogue in the listwise approach. Considering the quality of the pseudo answer, we design a note, i.e., “The reference answer may not be the correct answer, but it provides a pattern of the correct answer.”, adding before judging. We do experiments on listwise and pointwise respectively, and find that listwise is not sensitive to this note, note has a great influence on pointwise, as shown in Table 6 Finally, we choose prompt as shown in Figure 9 and Figure 10.
| listwise | pointwise | |||
| Note | Note | Note | Note | |
| =1 | 53.56 | 53.91 | 49.44 | 48.34 |
| =2 | 55.30 | 55.65 | 51.09 | 49.44 |
| =3 | 54.57 | 54.86 | 51.74 | 50.02 |
| =4 | 53.87 | 55.02 | 51.90 | 50.30 |
| =5 | 53.99 | 54.86 | 51.65 | 50.60 |
Moreover, for the implicit answer, we consider “what information is necessary to answer the question” and “which information is necessary to answer the question” in the prompt. Experiments are shown in Table 7. Finally, we choose prompt as shown in Figure 15. For the experiments of ChatGPT API, we have the cost of approximately $248.
| TREC | WebAP | |||
| which | what | which | what | |
| =1 | 49.39 | 48.34 | 26.06 | 25.81 |
| =2 | 50.72 | 50.13 | 28.67 | 25.38 |
| =3 | 52.05 | 50.94 | 28.36 | 25.04 |
| =4 | 51.71 | 51.54 | 25.46 | 26.87 |
| =5 | 51.54 | 50.66 | 27.35 | 26.69 |
A.2 Instruction of the Ranking Approach
For RankGPT, we directly use the instruction of Sun et al. (2023) for relevance ranking, as shown in Figure 13. For the relevance ranking in our ITEM, the instructions are shown in Figure 11 and Figure 12.
A.3 Instruction of Answer Generation
Li et al. (2023) utilize LLM to generate the missing information in the provided documents for the current question and then re-retrieve it as relevant feedback. Therefore, we have also designed two kinds of pseudo answers for utility judgments, i.e., (i) the explicit answer, which produces an answer based on the given information, and (ii) the implicit answer, which does not answer the question directly but gives the information necessary to answer the question. For “words/sentences”, the experiments found that the length of the explicit answers generated by Mistral using “sentences” is too long for factual questions, whereas Llama 3 and chatGPT use “sentences“ to generate answers of moderate length. So we only use “words“ on the TREC dataset and the NQ dataset using Mistral. The two instructions are shown in Figure 14 and Figure 15.
Appendix B Experiment Details
We utilized AI models such as ChatGPT to perform grammar corrections on portions of our paper.
B.1 Effect of Iteration Numbers
The performance of different LLMs in different datasets is shown in Table 8, Table 9, Table 10, Table 11 and Table 12.
| Method | TREC | WebAP | ||||||||||
| listwise | pointwise | listwise | pointwise | |||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Vanilla | 36.82 | 60.13 | 45.67 | 29.92 | 91.61 | 45.11 | 13.07 | 50.83 | 20.79 | 13.30 | 86.29 | 23.05 |
| UJ-ExpA | 48.51 | 61.15 | 54.10 | 28.12 | 96.27 | 43.53 | 18.83 | 54.16 | 27.94 | 14.65 | 91.82 | 25.27 |
| UJ-ImpA | 40.16 | 60.53 | 48.29 | 33.95 | 83.76 | 48.31 | 16.46 | 52.45 | 25.06 | 17.56 | 73.55 | 28.35 |
| 5-sampling | 46.64 | 59.56 | 52.31 | - | - | - | 20.61 | 56.22 | 30.16 | - | - | - |
| ITEM-As ExpA (=1) | 48.07 | 61.04 | 53.78 | 34.21 | 89.11 | 49.44 | 20.57 | 53.81 | 29.76 | 17.86 | 78.41 | 29.10 |
| ITEM-As ExpA (=2) | 50.58 | 61.86 | 55.65 | 35.87 | 88.73 | 51.09 | 21.11 | 50.85 | 29.83 | 18.27 | 82.00 | 29.88 |
| ITEM-As ExpA (=3) | 50.61 | 59.88 | 54.86 | 36.23 | 90.46 | 51.74 | 23.57 | 48.14 | 31.65 | 18.73 | 81.96 | 30.50 |
| ITEM-As ExpA (=4) | 50.01 | 61.15 | 55.02 | 36.41 | 90.36 | 51.90 | 21.44 | 44.62 | 28.96 | 19.19 | 80.59 | 31.00 |
| ITEM-As ExpA (=5) | 50.61 | 59.88 | 54.86 | 36.14 | 90.46 | 51.65 | 24.07 | 47.09 | 31.86 | 19.17 | 78.94 | 30.85 |
| ITEM-As ImpA (=1) | 39.97 | 64.62 | 49.39 | 30.98 | 89.38 | 46.01 | 16.88 | 57.13 | 26.06 | 17.10 | 81.65 | 28.28 |
| ITEM-As ImpA (=2) | 43.14 | 61.52 | 50.72 | 30.90 | 87.00 | 45.60 | 19.41 | 54.82 | 28.67 | 18.88 | 78.06 | 30.40 |
| ITEM-As ImpA (=3) | 44.43 | 62.82 | 52.05 | 31.68 | 87.99 | 46.59 | 19.21 | 54.20 | 28.36 | 18.69 | 77.77 | 30.13 |
| ITEM-As ImpA (=4) | 44.72 | 61.29 | 51.71 | 31.66 | 87.40 | 46.49 | 17.44 | 47.11 | 25.46 | 18.95 | 78.06 | 30.50 |
| ITEM-As ImpA (=5) | 44.63 | 60.98 | 51.54 | 31.80 | 89.32 | 46.91 | 18.98 | 48.88 | 27.35 | 19.05 | 76.69 | 30.52 |
| ITEM-ARs (=1) | 43.65 | 65.34 | 52.34 | - | - | - | 25.04 | 60.99 | 35.50 | - | - | - |
| ITEM-ARs (=2) | 45.10 | 65.46 | 53.40 | - | - | - | 24.42 | 51.97 | 33.23 | - | - | - |
| ITEM-ARs (=3) | 49.07 | 65.96 | 56.27 | - | - | - | 27.70 | 55.95 | 37.06 | - | - | - |
| ITEM-ARs (=4) | 50.96 | 62.32 | 56.07 | - | - | - | 23.77 | 53.40 | 32.90 | - | - | - |
| ITEM-ARs (=5) | 53.01 | 63.60 | 57.82 | - | - | - | 25.85 | 47.56 | 33.50 | - | - | - |
| Method | TREC | WebAP | ||||||||||
| listwise | pointwise | listwise | pointwise | |||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Vanilla | 34.67 | 85.80 | 49.39 | 31.42 | 98.47 | 47.64 | 12.69 | 77.15 | 21.79 | 14.65 | 87.36 | 25.09 |
| UJ-ExpA | 39.21 | 80.98 | 52.83 | 38.27 | 90.15 | 53.73 | 16.32 | 77.92 | 26.99 | 18.04 | 77.15 | 29.25 |
| UJ-ImpA | 33.92 | 83.36 | 48.22 | 38.68 | 71.47 | 50.20 | 15.57 | 82.79 | 26.22 | 17.22 | 47.61 | 25.29 |
| 5-sampling | 39.04 | 80.98 | 52.68 | - | - | - | 17.52 | 83.49 | 28.97 | - | - | - |
| ITEM-As ExpA (=1) | 39.68 | 82.88 | 53.66 | 37.58 | 84.84 | 52.09 | 17.54 | 63.67 | 27.50 | 19.65 | 74.31 | 31.08 |
| ITEM-As ExpA (=2) | 42.35 | 84.77 | 56.48 | 38.25 | 84.58 | 52.68 | 17.39 | 60.25 | 26.99 | 20.23 | 73.01 | 31.68 |
| ITEM-As ExpA (=3) | 42.00 | 84.15 | 56.03 | 37.84 | 85.50 | 52.46 | 19.12 | 62.87 | 29.32 | 20.91 | 74.63 | 32.67 |
| ITEM-As ExpA (=4) | 41.85 | 84.41 | 55.96 | 38.12 | 85.16 | 52.67 | 17.53 | 61.85 | 27.31 | 20.44 | 73.83 | 32.02 |
| ITEM-As ExpA (=5) | 42.36 | 84.15 | 56.35 | 37.35 | 84.69 | 51.84 | 18.94 | 62.87 | 29.12 | 20.88 | 75.45 | 32.71 |
| ITEM-As ImpA (=1) | 39.63 | 83.42 | 53.73 | 39.70 | 82.87 | 53.68 | 15.48 | 73.66 | 25.59 | 20.04 | 64.06 | 30.53 |
| ITEM-As ImpA (=2) | 38.75 | 85.63 | 53.35 | 38.15 | 82.36 | 52.14 | 15.50 | 76.47 | 25.77 | 18.54 | 62.69 | 28.62 |
| ITEM-As ImpA (=3) | 40.84 | 84.86 | 55.14 | 40.58 | 79.64 | 53.76 | 15.99 | 70.99 | 26.10 | 19.54 | 61.32 | 29.64 |
| ITEM-As ImpA (=4) | 38.88 | 82.74 | 52.90 | 39.34 | 81.74 | 53.12 | 15.03 | 74.41 | 25.01 | 19.72 | 59.95 | 29.68 |
| ITEM-As ImpA (=5) | 41.26 | 84.61 | 55.47 | 40.92 | 82.14 | 54.63 | 15.49 | 68.93 | 25.29 | 19.84 | 57.21 | 29.46 |
| ITEM-ARs (=1) | 34.53 | 84.17 | 48.97 | - | - | - | 20.05 | 72.88 | 31.44 | - | - | - |
| ITEM-ARs (=2) | 36.27 | 83.19 | 50.51 | - | - | - | 15.92 | 79.01 | 26.50 | - | - | - |
| ITEM-ARs (=3) | 38.04 | 82.68 | 52.10 | - | - | - | 17.93 | 76.87 | 29.08 | - | - | - |
| ITEM-ARs (=4) | 37.28 | 83.70 | 51.58 | - | - | - | 16.60 | 78.81 | 27.42 | - | - | - |
| ITEM-ARs (=5) | 40.25 | 81.37 | 53.86 | - | - | - | 17.04 | 74.83 | 27.75 | - | - | - |
| Method | TREC | WebAP | ||||||||||
| listwise | pointwise | listwise | pointwise | |||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Vanilla | 42.13 | 79.98 | 55.19 | 33.86 | 94.40 | 49.84 | 17.13 | 83.45 | 28.43 | 15.80 | 89.42 | 26.85 |
| UJ-ExpA | 45.74 | 77.36 | 57.49 | 32.06 | 96.19 | 48.09 | 19.51 | 69.86 | 30.50 | 16.23 | 88.74 | 27.44 |
| UJ-ImpA | 44.19 | 77.11 | 56.18 | 33.45 | 90.36 | 48.83 | 18.37 | 80.14 | 29.89 | 15.58 | 84.51 | 26.32 |
| 5-sampling | 50.78 | 74.77 | 60.49 | - | - | - | 20.70 | 65.83 | 31.49 | - | - | - |
| ITEM-As ExpA (=1) | 55.55 | 71.48 | 62.52 | 37.83 | 91.94 | 53.61 | 26.74 | 59.45 | 36.89 | 19.73 | 84.95 | 32.02 |
| ITEM-As ExpA (=2) | 57.95 | 70.40 | 63.57 | 40.74 | 93.04 | 56.67 | 29.43 | 60.58 | 39.62 | 19.62 | 78.62 | 31.40 |
| ITEM-As ExpA (=3) | 58.36 | 68.88 | 63.18 | 40.00 | 91.88 | 55.74 | 29.30 | 60.91 | 39.57 | 19.80 | 76.20 | 31.43 |
| ITEM-As ExpA (=4) | 58.48 | 70.67 | 64.00 | 40.25 | 93.38 | 56.25 | 29.11 | 61.03 | 39.42 | 20.48 | 79.63 | 32.58 |
| ITEM-As ExpA (=5) | 58.34 | 69.69 | 63.51 | 39.29 | 92.16 | 55.09 | 29.76 | 60.68 | 39.93 | 20.58 | 80.42 | 32.77 |
| ITEM-As ImpA (=1) | 54.36 | 65.08 | 59.24 | 40.89 | 82.20 | 54.61 | 24.79 | 64.37 | 35.80 | 18.78 | 67.00 | 29.34 |
| ITEM-As ImpA (=2) | 55.88 | 63.11 | 59.27 | 43.32 | 83.13 | 56.96 | 27.68 | 62.03 | 38.28 | 20.70 | 70.54 | 32.00 |
| ITEM-As ImpA (=3) | 57.33 | 64.17 | 60.56 | 41.66 | 80.48 | 54.90 | 30.01 | 63.60 | 40.78 | 21.51 | 66.77 | 32.54 |
| ITEM-As ImpA (=4) | 55.98 | 62.24 | 58.95 | 42.34 | 80.65 | 55.53 | 28.43 | 60.11 | 38.60 | 20.60 | 65.63 | 31.36 |
| ITEM-As ImpA (=5) | 56.63 | 62.19 | 59.28 | 41.49 | 83.57 | 55.45 | 29.05 | 60.66 | 39.29 | 21.51 | 68.03 | 32.68 |
| ITEM-ARs (=1) | 51.94 | 76.90 | 62.00 | - | - | - | 25.32 | 65.84 | 36.58 | - | - | - |
| ITEM-ARs (=2) | 53.77 | 76.19 | 63.05 | - | - | - | 25.55 | 59.26 | 35.70 | - | - | - |
| ITEM-ARs (=3) | 52.41 | 74.04 | 61.37 | - | - | - | 27.61 | 63.96 | 38.58 | - | - | - |
| ITEM-ARs (=4) | 52.75 | 73.78 | 61.52 | - | - | - | 28.84 | 61.85 | 39.34 | - | - | - |
| ITEM-ARs (=5) | 52.77 | 76.28 | 62.39 | - | - | - | 28.76 | 62.54 | 39.40 | - | - | - |
| References of RAG | Mistral | Llama 3 | ChatGPT | |||
| EM | F1 | EM | F1 | EM | F1 | |
| Golden Evidence | 46.09 | 62.59 | 64.45 | 76.64 | 66.40 | 76.86 |
| RocketQAv2 | 31.58 | 47.69 | 50.96 | 62.01 | 46.54 | 57.00 |
| Vanilla | 31.16 | 47.43 | 49.09 | 60.56 | 48.52 | 58.64 |
| UJ-ExpA | 32.76 | 48.46 | 49.63 | 61.10 | 47.72 | 58.01 |
| UJ-ImpA | 30.67 | 46.83 | 48.88 | 60.26 | 49.01 | 59.30 |
| 5-sampling | 33.24 | 48.84 | 48.72 | 60.71 | 48.90 | 58.97 |
| ITEM-As ExpA (=1) | 32.98 | 49.00 | 50.16 | 61.88 | 49.38 | 59.78 |
| ITEM-As ExpA (=2) | 34.31 | 50.08 | 50.48 | 62.32 | 49.22 | 59.99 |
| ITEM-As ExpA (=3) | 33.73 | 49.63 | 50.27 | 62.09 | 49.69 | 60.18 |
| ITEM-As ExpA (=4) | 34.21 | 50.07 | 50.43 | 62.20 | - | - |
| ITEM-As ExpA (=5) | 33.78 | 49.63 | 50.27 | 62.07 | - | - |
| ITEM-As ImpA (=1) | 32.17 | 48.51 | 50.37 | 61.89 | 48.75 | 58.99 |
| ITEM-As ImpA (=2) | 32.49 | 48.67 | 49.63 | 61.16 | 49.11 | 59.14 |
| ITEM-As ImpA (=3) | 32.39 | 48.47 | 49.68 | 61.48 | 48.69 | 58.94 |
| ITEM-As ImpA (=4) | 32.71 | 48.84 | 49.41 | 61.03 | - | - |
| ITEM-As ImpA (=5) | 32.33 | 48.44 | 49.73 | 61.42 | - | - |
| ITEM-ARs (=1) | 33.30 | 49.26 | 50.27 | 61.69 | 49.52 | 59.64 |
| ITEM-ARs (=2) | 33.57 | 49.16 | 50.70 | 61.92 | 49.01 | 59.75 |
| ITEM-ARs (=3) | 33.40 | 49.27 | 49.36 | 60.97 | 49.06 | 59.67 |
| ITEM-ARs (=4) | 33.46 | 49.24 | 49.84 | 61.54 | - | - |
| ITEM-ARs (=5) | 33.89 | 49.58 | 49.20 | 60.84 | - | - |
| NDCG@5 | NDCG@10 | NDCG@20 | Utility-judgments F1 | |
| 1 | 71.29 / 70.57 | 72.90 / 72.69 | 84.56 / 84.08 | 52.34 / 42.02 |
| 2 | 72.27 / 72.86 | 75.16 / 75.48 | 85.76 / 86.09 | 53.40 / 42.10 |
| 3 | 73.24 / 74.27 | 75.53 / 75.78 | 86.59 / 86.80 | 56.27 / 44.18 |
| 4 | 73.58 / 75.35 | 75.52 / 76.83 | 86.46 / 87.23 | 56.07 / 44.68 |
| 5 | 73.12 / 74.61 | 74.75 / 76.20 | 85.95 / 86.82 | 57.82 / 44.25 |
B.2 Iteration Stop Conditions
We consider two stop conditions for ITEM, i.e., utility judgments and answer. The different performances of Mistral with different stop conditions are in Table 13.
| Iteration stop condiction | TREC | WebAP | ||
| Listwise | Pointwise | Listwise | Pointwise | |
| Utility (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Utility (=2) | 55.65 | 51.09 | 29.83 | 29.88 |
| Utility (=3) | 54.86 | 51.74 | 31.65 | 30.50 |
| Utility (=4) | 55.02 | 51.90 | 28.96 | 31.00 |
| Utility (=5) | 54.86 | 51.65 | 31.86 | 30.85 |
| Answer 0.9 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.9 (=2) | 55.65 | 50.45 | 29.83 | 29.88 |
| Answer 0.9 (=3) | 54.90 | 50.88 | 31.65 | 30.50 |
| Answer 0.9 (=4) | 55.02 | 51.23 | 28.96 | 31.00 |
| Answer 0.9 (=5) | 54.90 | 50.79 | 31.86 | 30.85 |
| Answer 0.8 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.8 (=2) | 55.16 | 50.46 | 29.83 | 30.20 |
| Answer 0.8 (=3) | 54.35 | 50.85 | 31.65 | 30.81 |
| Answer 0.8 (=4) | 54.53 | 51.20 | 28.96 | 31.30 |
| Answer 0.8 (=5) | 54.35 | 50.76 | 31.86 | 31.35 |
| Answer 0.7 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.7 (=2) | 55.37 | 50.00 | 29.62 | 29.95 |
| Answer 0.7 (=3) | 54.93 | 50.31 | 31.58 | 30.75 |
| Answer 0.7 (=4) | 55.11 | 50.51 | 29.20 | 31.11 |
| Answer 0.7 (=5) | 54.93 | 50.28 | 31.77 | 31.31 |
| Answer 0.6 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.6 (=2) | 54.98 | 50.02 | 29.76 | 30.02 |
| Answer 0.6 (=3) | 55.34 | 50.27 | 32.28 | 30.55 |
| Answer 0.6 (=4) | 54.78 | 50.53 | 30.29 | 30.92 |
| Answer 0.6 (=5) | 55.34 | 50.23 | 32.14 | 31.05 |
| Answer 0.5 (=1) | 53.78 | 49.44 | 29.76 | 29.10 |
| Answer 0.5 (=2) | 55.30 | 49.98 | 31.42 | 30.15 |
| Answer 0.5 (=3) | 55.39 | 50.23 | 32.97 | 30.17 |
| Answer 0.5 (=4) | 55.10 | 50.42 | 31.73 | 29.92 |
| Answer 0.5 (=5) | 55.39 | 50.16 | 32.50 | 29.92 |
B.3 Quality of Utility Judgments
The retrieved data we use are a four-point scale, and we consider the highest level as having utility. To see the different points in our results for utility judgments, we analyzed the various ranks as a percentage of the results on TREC for mistral in Table 14.
| label1 | label2 | label3 | |
| =1 | 82.08 | 68.34 | 48.07 |
| =2 | 83.86 | 69.53 | 50.58 |
| =3 | 84.23 | 71.06 | 50.61 |
| =4 | 84.63 | 70.18 | 50.01 |
| =5 | 84.52 | 70.69 | 50.61 |
From Table 14, we can see that almost 70% of the results of utility judgments are very relevant to the question.
B.4 values in ITEM-Ar
Different ranking performance of values in ITEM-Ar is shown in Table 15. Considering the performance of utility ranking and utility judgments, we set =5.
| , | Ranking | Utility judgments | ||||||
| N@1 | N@3 | N@5 | N@10 | N@20 | P | R | F1 | |
| =1, =1 | 72.76 | 71.27 | 70.57 | 72.69 | 84.08 | 53.66 | 24.09 | 33.25 |
| =1, =2 | 76.02 | 71.54 | 71.38 | 73.66 | 84.78 | 58.54 | 28.73 | 38.54 |
| =1, =3 | 77.24 | 72.83 | 71.83 | 73.87 | 85.20 | 59.76 | 28.84 | 38.90 |
| =1, =4 | 77.24 | 73.04 | 71.91 | 73.90 | 85.25 | 59.76 | 28.84 | 38.90 |
| =1, =5 | 76.02 | 72.11 | 71.42 | 73.45 | 84.98 | 58.54 | 28.71 | 38.53 |
| =5, =1 | 72.76 | 71.27 | 70.57 | 72.69 | 84.08 | 33.17 | 57.31 | 42.02 |
| =5, =2 | 78.46 | 73.74 | 72.86 | 75.48 | 86.09 | 32.93 | 58.37 | 42.10 |
| =5, =3 | 79.27 | 75.00 | 74.27 | 75.78 | 86.80 | 34.15 | 62.57 | 44.18 |
| =5, =4 | 79.67 | 75.92 | 75.35 | 76.83 | 87.23 | 35.12 | 61.40 | 44.68 |
| =5, =5 | 79.67 | 75.32 | 74.61 | 76.20 | 86.82 | 34.63 | 61.25 | 44.25 |
| =10, =1 | 72.76 | 71.27 | 70.57 | 72.69 | 84.08 | 22.56 | 68.03 | 33.88 |
| =10, =2 | 78.05 | 72.64 | 72.90 | 75.48 | 85.74 | 23.66 | 75.47 | 36.02 |
| =10, =3 | 80.89 | 76.58 | 74.54 | 76.30 | 86.94 | 23.78 | 75.65 | 36.19 |
| =10, =4 | 78.05 | 74.70 | 72.85 | 75.12 | 85.72 | 24.51 | 74.17 | 36.85 |
| =10, =5 | 79.67 | 75.60 | 74.84 | 76.54 | 86.88 | 23.66 | 74.42 | 35.90 |
Appendix C Case Study
Appendix D Annotations in NQ Dataset
To evaluate the performance of utility judgments under different values of , we relabeled the retrieved passages following Stelmakh et al. (2022). Specifically, we used a QA model based on RoBERTa Liu et al. (2019) to extract answers from the retrieved passages. A passage is considered utility if the extracted answers match the ground-truth answers. The advantage of this labeling method is its relatively high accuracy, while its disadvantage is the potentially low recall rate.
Appendix E Datasets
Detailed statistics of the experimental datasets are shown in Table 16.
| Dataset | #Psg | #PsgLen | #Q | #Rels/Q |
| TREC | 8.8M | 93 | 82 | 212.8 |
| WebAP | 379k | 45 | 73 | 76.4 |
| NQ | 21M | 100 | 1868 | 1.0 |
Appendix F Answer Passage Retrieval
Non-factoid questions are usually expected longer answers, such as sentence-level or passages-level Keikha et al. (2014a); Yang et al. (2016); Keikha et al. (2014b). Yang et al. (2016) developed an annotated dataset for answer passage retrieval called WebAP, which has an average of 76.4 qrels per query. Cohen et al. (2018) and Hashemi et al. (2020) introduced the WikiPassageQA dataset and ANTIQUE dataset for answer passage retrieval, respectively. Compared to the WebAP dataset, WikiPassageQA and ANTIQUE have incomplete annotations, with an average of 1.7 qrels and 32.9 qrels per query Hashemi et al. (2019, 2020). Bi et al. (2019) created the PsgRobust dataset for answer passage retrieval, which is built on the TREC Robust collection Voorhees et al. (2003) following a similar approach to WebAP but without manual annotation.
Appendix G Parameters Comparison
The output of -sampling each time contains explicit answers and utility judgments. If the question length is , the total length of the input passages is , and the average length of a single passage is , then the k-sampling input cost is . If the average length of the output explicit answer is , and the average length of the output utility judgments is , then the k-sampling output cost is . Taking ITEM-As as an example, with a maximum of three iterations, the maximum input cost for utility judgments is . For answer generation, the longest input is and the shortest is . Therefore, the maximum input cost for ITEM-As is and the minimum is . The maximum output cost is . Therefore, in order to ensure fairness in the calculation of large language model parameters, we choose =5.