### v2

# An Iterative Utility Judgment Framework Inspired by Philosophical Relevance via LLMs

Abstract
Relevance and utility are two frequently used measures to evaluate the effectiveness of an information retrieval (IR) system. Relevance emphasizes the aboutness of a result to a query, while utility refers to the result’s usefulness or value to an information seeker. In Retrieval-Augmented Generation (RAG), high-utility results should be prioritized to feed to LLMs due to their limited input bandwidth. Re-examining RAG’s three core components—relevance ranking derived from retrieval models, utility judgments, and answer generation—aligns with Schutz’s philosophical system of relevances, which encompasses three types of relevance representing different levels of human cognition that enhance each other. These three RAG components also reflect three cognitive levels for LLMs in question-answering. Therefore, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to promote each step in RAG. We conducted extensive experiments on retrieval (TREC DL, WebAP), utility judgment task (GTI-NQ), and factoid question-answering (NQ) datasets. Experimental results demonstrate significant improvements of ITEM in utility judgments, ranking, and answer generation upon representative baselines111Our code and benchmark can be found at https://anonymous.4open.science/r/ITEM-B486/..
An Iterative Utility Judgment Framework Inspired by Philosophical Relevance via LLMs
Hengran Zhang1,2,3 Keping Bi1,2,3 Jiafeng Guo1,2, 3 Xueqi Cheng1,2,3 1Key Laboratory of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences; 2State Key Laboratory of AI Safety; 3University of Chinese Academy of Sciences; {zhanghengran22z, bikeping, guojiafeng, cxq}@ict.ac.cn,
1 Introduction
Relevance and utility are two frequently used measures to evaluate Information Retrieval (IR) performance Saracevic (1996, 1975); Saracevic et al. (1988). Relevance usually refers to topical relevance that measures the degree of fit between the subjects of a query and retrieved items, and the criteria of “aboutness” is used Saracevic et al. (1988); Schamber and Eisenberg (1988). In contrast, utility refers to the “usefulness” of retrieval items to an information seeker, of which the criterion is their “value” to the user Saracevic (1996); Saracevic et al. (1988). As the example from the TREC DL dataset shown on Figure 1, topical relevance does not necessarily mean utility, while utility indicates a higher standard of relevance. Since topical relevance is relatively easy to observe and measure Schamber et al. (1990), the studies of IR models have been primarily focused on improving relevance for a long time Bruce (1994).
In the modern LLM era, Retrieval-Augmented Generation (RAG) has become a hot research topic that equips LLMs with external knowledge Xie et al. (2023); Shi et al. (2023); Izacard et al. (2023); Su et al. (2024); Glass et al. (2022). Given the constrained bandwidth of LLM inputs, it is essential to prioritize high-value results to guide LLMs. Consequently, utility needs to be emphasized more than topical relevance in RAG. More recently, Zhang et al. (2024) highlighted the use of LLMs for utility judgments. In this paper, we aim to further promote the utility judgment performance of LLMs so that RAG can be enhanced by high-utility references.
Schutz’s Philosophical Theory of Relevance. Relevance is foundational in information retrieval (IR) and remains widely debated Mizzaro (1998). Saracevic (1996) discussed the nature of relevance in the IR system as the effectiveness of interactive exchange on different levels, and they are non-independent interdependencies, which are primarily influenced by Schutz’s philosophical theory of relevance. Schutz considered relevance as the property that determines the connections and relations in our lifeworld. He identified three types of basic and interdependent relevance that interact dynamically within a “system of relevancies” Saracevic (1996); Schutz (1970): (i) Topical relevance, which refers to the perception of what is separated from one’s experience to form one’s present object of concentration; (ii) Interpretational relevance, which involves the past experiences in understanding the currently concerned object; and (iii) Motivational relevance, which refers to the course of action to be adapted based on the interpretations. The motivational relevance, in turn, helps obtain additional materials to become a user’s new experience, which further facilitates topical and interpretational relevance. Schutz posited that one’s perception of the world may be enhanced by this dynamic interaction, as shown in Figure 2. By incorporating utility judgments into RAG, we can re-examine its three components: topical relevance or relevance ranking derived from retrieval models, utility judgments, and answer generation. Topical relevance is an emerging focus on a topic, utility is the deeper understanding of the topic, and answers indicate the final solution based on the interpretations and will guide users’ actions. Therefore, topic relevance, utility, and derived answer also reflect three cognitive levels for LLMs in question-answering, from low to high, i.e., aboutness, the value for deriving an answer, and the derived answer.
Iterative utiliTy judgmEnt fraMework (ITEM). Inspired by the philosophical theory of relevance, we believe the dynamic interactions between the three components in RAG can promote the performance of each step. To verify the idea, we leverage LLMs to perform each step in RAG shown in Figure 2, i.e., relevance ranking, utility judgments (classification), and answer generation. We propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps. ITEM has two variants depending on whether relevance ranking is involved in the iterations. We are curious to see which option will be better for the tasks: fewer iterations with more components in an iteration, more iterations with fewer components in an iteration, or more iterations with more components.
We experiment on various information-seeking tasks, i.e., multi-grade passage retrieval on TREC DL Craswell et al. (2020), multi-grade non-factoid answer passage retrieval on WebAP Yang et al. (2016), utility judgments benchmark on GTI-NQ Zhang et al. (2024), and factoid QA on NQ Kwiatkowski et al. (2019). For multi-grade passage retrieval, we consider the ones with the highest grade to be of utility and we focus on the performance of utility judgments and topical relevance ranking. For factoid QA, we emphasize the answer accuracy. Experimental results have demonstrated that ITEM can significantly outperform competitive baselines, including various single-shot judgment approaches in terms of utility judgments, ranking of topical relevance, and answer generation, which confirms the viability of adaptation of Schutz’s viewpoint of the relevance system into RAG. We also find that: 1) for difficult tasks (i.e., utility judgments of non-factoid answer passages in WebAP) and complicated candidate passage list (i.e., GTI-NQ), more components in the iteration and multiple iterations are usually more beneficial; 2) for multi-grade relevance ranking tasks, using utility as ranking criterion leads to significantly better multi-grade relevance performance than relevance ranking results even when utility judgment is involved in the iteration; 3) for factoid QA tasks, more iterations with fewer components performs the best, indicating that more components and more iterations are not always needed, especially for simpler tasks.
2 Related Work
Multi-dimensional relevance. The concept of “relevance” is central to information retrieval theory. Researchers have extensively debated its definition and measurement Mizzaro (1997). Early approaches primarily defined and assessed relevance through exact term matching Vickery (1959) or logical entailment Hillman (1964). However, subsequent empirical studies revealed the limitations of system-oriented relevance analysis, prompting diverse perspectives on relevance Saracevic (1975); Swanson (1986); Saracevic (1996); Lancaster (1968); Goffman and Newill (1964); Kemp (1974); Bruce (1994). For example, Cooper (1971) introduced logical relevance and utility. Saracevic (1996) summarized five frameworks for information science: systems, communication, situational, psychological, and interaction frameworks, and categorized five distinct types of relevance, i.e., 1) system or algorithmic relevance; 2) topical or subject relevance; 3) cognitive relevance or pertinence; 4) situational relevance or utility; and 5) motivational or affective relevance. Bruce (1994) explored cognitive dimensions of relevance. Over time, scholarly consensus has coalesced around two primary perspectives: the system view and user view, with topical relevance and utility serving as their respective representative frameworks.
Utility-Focused Information Retrieval. Utility is a distinct measure of relevance compared to topical relevance Zhao et al. (2024); Saracevic et al. (1988); Saracevic (1975, 1996); Ji et al. (2024); Zhang et al. (2023), and more recently, Zhang et al. (2024) highlighted the use of LLMs for utility judgments. However, Zhang et al. (2024) only conducted a preliminary exploration of LLMs in utility judgments. Our work aims to further explore how to improve the performance of utility judgments for LLMs.
Retrieval-Augmented Generation (RAG). RAG approaches are widely employed to mitigate the hallucination issues in large language models (LLMs) Xie et al. (2023); Zhou et al. (2024); Su et al. (2024). The current RAG approaches are categorized as follows: (i) single-round retrieval Borgeaud et al. (2022); Lewis et al. (2020); Glass et al. (2022); Izacard et al. (2023); Shi et al. (2023), which involves using the initial input as a query to retrieve information from an external corpus and then the information is incorporated as part of the input for the model; and (ii) multi-round retrieval Su et al. (2024); Jiang et al. (2023b); Ram et al. (2023); Khandelwal et al. (2020); Trivedi et al. (2023), which need multi-round retrieval based on feedback from LLMs.
Iterative Relevance Feedback via LLMs. Recent works Li et al. (2023); Shao et al. (2023) have achieved great success in using LLMs to obtain the information needs of the question as pseudo-relevance feedback for iterative retrieval. They posit that a single retrieval may not yield comprehensive information, thus requiring multiple retrievals. In contrast, our methodology involves making iterative utility judgments on the results obtained from a single retrieval. Given the substantial operational costs associated with retrieval systems, the expense incurred from conducting multiple retrievals for a single query becomes even more prohibitive.
3 Utility Judgments (UJ) via LLMs
Schutz’s philosophy of relevances encompasses three types of relevance: topical, interpretational, and motivational relevance, representing different levels of human cognition, and their dynamic interactions can enhance each other. By incorporating utility judgments into RAG and re-examining its three components, i.e., relevance ranking derived from the retrieval models, utility judgments, and answer generation, we realize they closely correspond to Schutz’s philosophical system of relevance. Topic relevance, utility, and derived answer also reflect three cognitive levels for LLMs in question-answering, from low to high, i.e., aboutness, the value for deriving an answer, and the derived answer. Inspired by Schutz’s theory, we presume they can also interact with each other and enhance each other. Therefore, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) for utility judgments.
3.1 Notations and Definitions
Given a question and a list of retrieved passages based on topical relevance, the goal of utility judgments for LLMs is to identify a set of passages , is the number of passage with utility judged by LLMs. There are two typical input approaches for LLMs to construct the set : pointwise and listwise: The pointwise approach independently evaluates the utility of individual passages, whereas the listwise method assesses the utility of multiple passages with the list input.
3.2 Single-Shot Utility Judgments
The most common approach to judge utility for the is to perform a single-shot utility judgment, i.e., , where is the instruction. Zhang et al. (2024) proposed to generate a pseudo-answer while conducting the utility judgment task, which can help LLMs to judge utility better, i.e., .
3.3 Iterative utiliTy judgmEnt fraMework (ITEM)
Inspired by Schutz’s theory of relevance in philosophy, we propose an Iterative Utility Judgment Framework (ITEM) for RAG. As illustrated in Figure 3, our framework dynamically iterates among topical relevance ranking, pseudo-answer generation, and utility judgments. We propose two types of loops where two or three components in RAG interact with each other iteratively (ITEM-A and ITEM-AR).
ITEM with Answering in the Loop (ITEM-A). Formally, at each iteration (), given the pseudo answer generated based on the utility judgment result from the previous iteration, we perform utility judgments on the candidate passages list to obtain a set of passages with utility :
where represents the answer prompts for LLMs (as detailed in Figure 4), can be in two forms (details are shown in the Figure 4): (i) explicit answerto the question ; (ii) implicit answerthat specifies the necessary information to answer the question . denotes the utility judgment prompts for LLMs (as detailed in Figure 5). We consider as the initial candidate set, where represents the initial results ranked by a retriever such as BM25 Robertson et al. (2009).
ITEM with both Answering and Ranking of Topical Relevance in the Loop (ITEM-AR). In the ITEM-A framework, topical relevance is not updated during the iteration process. To incorporate dynamic updating of topical relevance, we integrate a relevance ranking task into the ITEM framework, ensuring that all three tasks are executed in a loop. Formally, at iteration (), the answer is generated based on the judging result from the previous iteration. Subsequently, given , the passage list from the previous iteration is ranked based on the relevance to the question, yielding a new ranked list . Finally, the judging result is derived using the ranked list and the answer :
where is the relevance ranking prompt for LLMs (as detailed in Figure 6), respectively.
Overall. At iteration , we have two ways to produce the set : (i) Set-based approach: Asking LLMs to identify the set of passages that have utility using listwise and pointwise input forms, which called ITEM-As or ITEM-ARs variants; (ii) Rank-based approach: Requesting LLMs to provide a ranked passage list based on utility (utility ranking prompt is shown in Appendix H.2) using the listwise input approach and considering the top- passages in the list to build , which called ITEM-Ar or ITEM-ARr variants. We set and more details of are shown in Appendix A.3. We find that ITEM-ARr does not improve ranking performance as well as ITEM-Ar (see Appendix A.4 for experimental analysis), so we do not employ ITEM-ARr in the ranking experiment. The rank-based approach has poor performance on the utility judgment task (details can be found in Appendix A.4), so we only show the performance of the set-based approach on the utility judgment task.
We stop the iteration when at most (=3 in our paper) iterations are reached or the set of selected passages does not change, i.e., or . Full details of all prompts can be found in Appendix H.
4 Experimental Setup
4.1 Datasets
Our experiments are conducted on four benchmark datasets, including two retrieval datasets, i.e., TREC DL Craswell et al. (2020) and WebAP Yang et al. (2016), a utility judgment dataset, i.e., GTI-NQ Zhang et al. (2024), and an open-domain question answer (ODQA) dataset, i.e., NQ Kwiatkowski et al. (2019). Detailed statistics of the experimental datasets are shown in Table Appendix C. We use two representative retrievers to gather candidate passages in for utility judgments on TREC DL, WebAP, and NQ datasets. Construction details can be found in Appendix F.
TREC DL. We use the TREC-DL19 and TREC-DL20 datasets Craswell et al. (2020). Judgments of TREC DL are on a four-point scale, i.e., “perfectly relevant”, “highly relevant”, “related”, and “irrelevant”. We consider the passages that are “perfectly relevant” to have utility. We filter questions of two datasets that contain the passages labeled “perfectly relevant’ and combine them to form a whole dataset, i.e., the TREC DL.
WebAP. WebAP Yang et al. (2016) is a non-factoid answer passage collection built on Gov2. Non-factoid questions usually require longer answers, such as sentence-level or passage-level Keikha et al. (2014a); Yang et al. (2016); Keikha et al. (2014b). Relevant passages are annotated and categorized as “perfect”, “excel”, “good”, and “fair”. Similarly to TREC DL, we considered the “perfect” passages to have utility.
NQ. Natural Questions (NQ) consist of factoid questions issued to the Google search engine Kwiatkowski et al. (2019). Each question is annotated with a long answer (typically a paragraph) and a short answer (one or more entities). Following Zhang et al. (2024), we use the questions that have long answers in our experiments.
GTI-NQ. Ground-truth inclusion (GTI) benchmark is constructed by Zhang et al. (2024) for utility judgment task. The GTI-NQ constructs a candidate passage set of 10 passages for each query sourced from the NQ dataset, comprising the long answer (designated as the utility passage), highly relevant noisy passages, weakly relevant noisy passages, and counterfactual passages.
4.2 Evaluation metrics
For the utility judgments task, we evaluate the results of judgments using Precision, Recall, and micro F1. For the ranking task, we use the normalized discounted cumulative gain (NDCG) Järvelin and Kekäläinen (2017) to evaluate the ranking performance. For the answer generation task, we use the standard exact match (EM) metric and F1.
4.3 LLMs
We conduct our experiments using several representative LLMs, i.e., (i) ChatGPT OpenAI (2022) (we use the gpt-3.5-turbo-1106 version), (ii) Mistral Jiang et al. (2023a) (the Mistral-7B-Instruct-v0.2 version), and (iii) Llama 3 Meta (2024) (the Meta-Llama-3-8B-Instruct version). To ensure the reproducibility of the experiments, the temperature for all experiments is set to 0.
| Method | WebAP | TREC DL | ||||||||||
| Listwise | Pointwise | Listwise | Pointwise | |||||||||
| M | L | C | M | L | C | M | L | C | M | L | C | |
| Vanilla | 20.79 | 21.79 | 28.43 | 23.05 | 25.09 | 26.85 | 45.67 | 49.39 | 55.19 | 45.11 | 47.64 | 49.84 |
| UJ-ExpA | 27.94 | 26.99 | 30.50 | 25.27 | 29.25 | 27.44 | 54.10 | 52.83 | 57.49 | 43.53 | 53.73 | 48.09 |
| UJ-ImpA | 25.06 | 26.22 | 29.89 | 28.35 | 25.29 | 26.32 | 48.29 | 48.22 | 56.18 | 48.31 | 50.20 | 48.83 |
| -sampling | 30.16 | 28.97 | 31.49 | - | - | - | 52.31 | 52.68 | 60.49 | - | - | - |
| ITEM-As ExpA (1) | 29.76 | 27.50 | 36.89 | 29.10 | 31.08 | 32.02 | 53.78 | 53.66 | 62.52 | 49.44 | 52.09 | 53.61 |
| ITEM-As ImpA (1) | 26.06 | 25.59 | 34.97 | 28.28 | 30.53 | 29.34 | 49.39 | 53.73 | 58.11 | 46.01 | 53.68 | 54.61 |
| ITEM-ARs ExpA(1) | 35.50 | 31.44 | 36.58 | - | - | - | 52.34 | 48.97 | 62.00 | - | - | - |
| ITEM-As ExpA (3) | 31.65 | 29.32 | 39.57 | 30.50 | 32.67 | 31.43 | 54.86 | 56.03 | 63.18 | 51.74 | 52.46 | 55.74 |
| ITEM-As ImpA (3) | 28.36 | 26.10 | 40.78 | 30.13 | 29.64 | 32.54 | 52.05 | 55.14 | 60.56 | 46.59 | 53.76 | 54.90 |
| ITEM-ARs ExpA(3) | 37.06 | 29.08 | 38.58 | - | - | - | 56.27 | 52.10 | 61.37 | - | - | - |
4.4 Baselines
We utilize the following baselines on the utility judgments task and question answering performance based on the utility judgment results:
Single-shot utility judgments. (i) Vanilla: Ask LLMs to provide utility judgments based on the instruction directly. (ii) UJ-ExpA: Utility judgments and provide explicit answers simultaneously through a single output, which is shown to be effective in Zhang et al. (2024). (iii) UJ-ImpA: Utility judgments and provide implicit answers that are necessary to answer the question through a single output.
-sampling. Zhang et al. (2024) proposed -sampling to alleviate the sensitivity of LLMs to input order. Specifically, the -sampling method randomizes the order of the input passage list times in addition to the original input and aggregates the utility judgment results through voting. For fair comparison, we use the , more details are on Appendix E.
To evaluate the effectiveness of the proposed ITEM framework in ranking tasks, we are using a verberlized ranking. Therefore, we also employ another verberlized ranking method, i.e., RankGPT Sun et al. (2023) as our main baseline, which uses the LLMs to directly rank input passages based on their relevance to the query.
| Method | Llama3 8B | ChatGPT | ||
| Listwise | Pointwise | Listwise | Pointwise | |
| Vanilla | 43.38 | 28.55 | 59.37 | 35.31 |
| UJ-ExapA | 47.07 | 39.32 | 66.13 | 37.17 |
| UJ-ImpA | 43.31 | 38.72 | 57.40 | 37.29 |
| k-sampling | 49.20 | - | 71.17 | - |
| ITEM-As-ExpA(1) | 49.26 | 47.52 | 72.44 | 54.89 |
| ITEM-As-Imp(1) | 47.47 | 37.98 | 68.92 | 43.17 |
| ITEM-ARs-ExpA(1) | 50.77 | - | 74.43 | - |
| ITEM-As-ExpA(3) | 49.73 | 48.90 | 73.55 | 55.45 |
| ITEM-As-Imp(3) | 48.03 | 38.34 | 69.68 | 43.58 |
| ITEM-ARs-ExpA(3) | 51.22 | - | 76.34 | - |
5 Experimental Results of LLMs
This section will present the performance of each task within our ITEM framework. By default, the pseudo answer is the explicit answer in all experiments, if not specified otherwise.
5.1 Utility Judgments Results
Table 1 shows the micro F1 performance on the TREC DL and WebAP datasets using three LLMs. Further, we utilize a better-performing open-source LLM, i.e., Llama-3.1-8B, and a closed LLM, i.e., ChatGPT, to conduct experiments on GTI-NQ, as shown in Table 2. Since ITEM-Ar and ITEM-ARr have poor F1 performance in utility judgments (refer to Table 12 in Appendix A.4 for details), we restrict our experiments to ITEM-As and ITEM-ARs in this section.
ITEM with a Single Iteration vs. Baselines. All LLMs using our ITEM with a single iteration generally outperform the single-shot utility judgments on three datasets and may even surpass the -sampling method. For example, ChatGPT on the TREC DL dataset using our ITEM-As ExpA and ImpA in the listwise approach improve the F1 performance by 8.7% and 3.4% over UJ-ExpA and UJ-ImpA, respectively. Explicit generation of pseudo-answers by LLMs enhances their performance in utility judgment tasks, highlighting the importance of task interaction. Moreover, concurrent execution of answer generation and utility judgment within a single inference cycle yields inferior performance compared to sequential task execution through separate reasoning phases.
ITEM with Multiple Iterations vs. ITEM with Single Iteration. All LLMs using our ITEM-A and ITEM-AR generally demonstrate improved performance with multiple iterations compared to single iterations on all three datasets. For instance, on the WebAP dataset, Mistral, Llama 3, and ChatGPT (using our ITEM-A ExpA) improved their F1 scores in the listwise approach by 6.4%, 6.6%, and 7.3%, respectively, after multiple iterations. Moreover, our method achieves state-of-the-art performance compared to all baselines by leveraging the iterative framework. The performance improvement from multiple iterations underscores the significance of iterative interaction and further supports Schutz’s interactive framework. However, in some specific cases, multiple iterations may not outperform single iterations, likely due to the unpredictable nature of zero-shot settings. ChatGPT outperforms other LLMs on all datasets using both input approaches.
ITEM-As vs. ITEM-ARs. In our utility-emphasized iterative RAG framework, ITEM-As and ITEM-ARs are the two major methods we propose. From Table 1, we find that ITEM-ARs works better than ITEM-As most of the time for complex questions (WebAP, all the questions are non-factoid) and the complex candidate passage list (GTI-NQ, containing different types of passage), indicating complicated question or passage list need more components in the loop. For TREC DL, which contains factoid questions, we find that ITEM-ARs is worse than ITEM-As most times on TREC DL. This is reasonable since factoid questions are relatively easier to answer and may not need more components involved in the iteration.
Listwise vs. Pointwise. The general performance of utility judgments for LLMs is better with the listwise approach than with the pointwise approach. The primary rationale lies in the listwise approach exposes the LLM to broader contextual information, thereby facilitating more effective interaction during the LLMs in judging the passages’ utility.
| Method | Mistral | Llama 3 | ChatGPT | |||
| TREC | WebAP | TREC | WebAP | TREC | WebAP | |
| 58.69 | 21.89 | 58.69 | 21.89 | 58.69 | 21.89 | |
| RankGPT | 69.81 | 29.34 | 75.61 | 41.73 | 80.56 | 42.49 |
| ITEM-Ar (1) | 70.57 | 37.11 | 73.95 | 40.89 | 80.79 | 50.30 |
| ITEM-ARs (1) | 71.29 | 37.48 | 77.22 | 43.80 | 81.38 | 48.42 |
| ITEM-Ar (3) | 74.27 | 43.80 | 77.34 | 45.88 | 83.12 | 51.61 |
| ITEM-ARs (3) | 73.24 | 45.45 | 74.80 | 44.87 | 82.89 | 48.80 |
| Method | Llama 3 | ChatGPT | ||
| @5 | @10 | @5 | @10 | |
| 29.46 | 45.26 | 29.46 | 45.26 | |
| RankGPT | 71.50 | 74.05 | 77.27 | 78.64 |
| ITEM-Ar(m=1) | 74.36 | 76.91 | 85.99 | 87.26 |
| ITEM-ARs(m=1) | 75.46 | 77.75 | 84.54 | 85.14 |
| ITEM-Ar(m=3) | 75.95 | 78.18 | 87.48 | 88.47 |
| ITEM-ARs(m=3) | 76.38 | 78.56 | 85.95 | 86.39 |
5.2 Ranking Performance
We also assess whether the ranking performance has been improved within ITEM on retrieval datasets (Table 3) and utility judgments benchmark (Table 4). In terms of ranking performance, we consider two rankings: relevance ranking (ITEM-ARs) and utility ranking (ITEM-Ar). We can observe that: (i) Our ITEM with single iteration significantly improves the ranking of topical relevance performance compared to the RankGPT. For instance, relevance ranking outperforms RankGPT in NDCG@5 by 2.1% on the TREC dataset. The performance improvement may stem from the interaction between tasks. (ii) After iterations, relevance and utility ranking performance have been improved on all datasets and all LLMs. The ranking benefits from our dynamic iterative framework, confirming Schutz’s theory of dynamic iterative interaction. (iii) From Table 3&4, we can find that ITEM-ARs is generally better than ITEM-Ar when . However, when , it may have the opposite performance. The possible reason is that when is small, the answer is very good, and utility is more dependent on the answer than relevance. However, as iterations occur, the quality of the answer is better, and utility performance is gradually improved, while relevance does not have as obvious improvement effect as utility in iterations. This indicates that reranking based on utility performs better than including more components in the loop, but still using relevance as the ranking criterion. These findings further confirm the importance of the concept of utility in RAG.
5.3 Results of Answer Generation
In the answer generation task, the results of utility judgments are fed to LLMs for answer generation. We use the factoid QA dataset (i.e., NQ) for answer generation evaluation, as shown in Table 5. From Table 1&2, we find that the listwise approach generally outperforms the pointwise approach for utility judgments. Consequently, our answer generation experiments utilize only the listwise utility judgments.
| References | Mistral | Llama 3 | ChatGPT | |||
| EM | F1 | EM | F1 | EM | F1 | |
| Golden | 46.09 | 62.59 | 64.45 | 76.64 | 66.40 | 76.86 |
| 31.58 | 47.69 | 50.96 | 62.01 | 46.54 | 57.00 | |
| Vanilla | 31.16 | 47.43 | 49.09 | 60.56 | 48.52 | 58.64 |
| UJ-ExpA | 32.76 | 48.46 | 49.63 | 61.10 | 47.72 | 58.01 |
| UJ-ImpA | 30.67 | 46.83 | 48.88 | 60.26 | 49.01 | 59.30 |
| 5-sampling | 33.24 | 48.84 | 48.72 | 60.71 | 48.90 | 58.97 |
| ITEM-As (1) | 32.98 | 49.00 | 50.16 | 61.88 | 49.38 | 59.78 |
| ITEM-ARs (1) | 33.30 | 49.26 | 50.27 | 61.69 | 49.52 | 59.64 |
| ITEM-As (3) | 33.73 | 49.63 | 50.27 | 62.09 | 49.69 | 60.18 |
| ITEM-ARs (3) | 33.40 | 49.27 | 49.36 | 60.97 | 49.06 | 59.67 |
The following observations can be made from Table 5: (i) ITEM outperforms baselines on all metrics on all LLMs (except for the EM score of Llama 3), indicating that ITEM can help the LLMs to find better evidence for generating answers. (ii) Similar to Table 1&2, when the , ITEM-ARs performs better than ITEM-As, which shows the importance of relevance reranking in ITEM. However, as the number of iterations increases, ITEM-As performs better than ITEM-ARs. We are keen to discern the optimal choice for different tasks: 1) More components and more iterations are not always needed, especially for simpler tasks; 2) Fewer iterations with numerous components, or increased iterations with few components.
6 Further Analyses
Iteration Rounds. Figure 7 shows the performance of (a): utility judgments under ITEM-As and (b): ranking with varying maximum iteration rounds .
We observe the following: 1) Varying the value of affects the performance of utility judgments, ranking. 2) Based on empirical observations balancing the cost and performance, the was operationally configured with distinct values for different question types on utility judgments (=3 in our paper on all experiments for fair comparison): =2 for factoid questions, whereas =3 was implemented for non-factoid questions in practical applications. 3) Utility ranking generally outperforms relevance ranking, which confirms the effectiveness of utility in the ranking task.
Iteration Stop Conditions. In addition to utility judgments, we also consider the pseudo answer generation performance in ITEM as a stopping condition. Specifically, we calculate the ROUGE-L score Lin (2004) of the answer in two iterations and stop the iteration if the ROUGE-L of and is greater than . The utility judgments performance of different iteration stop conditions are shown in Figure 8. The results show that using different stopping conditions affects the performance of utility judgments. However, using the answer as a stopping condition, different LLMs on different datasets may need to look for different , which is not very flexible.
7 Inference Efficiency
Table 14 in Appendix G shows the inference efficiency analysis of our ITEM. Our iteration framework surpasses -sampling in computational efficiency during inference. The proposed approach demonstrates potential for large-scale retrieval data annotation, where ITEM-As emerges as an optimal solution by simultaneously enhancing both performance and operational efficiency, thereby facilitating utility annotation for retrieval systems.
8 Case Study
Figure 9 in Appendix B presents two cases from the TREC DL dataset using Mistral under ITEM-As. For the first question in Figure 9, the first pseudo-answer, though relatively correct, includes irrelevant information, leading to a misjudgment of “Passage-2” as “utility”. Based on the results of the first round of utility judgments, the second round of the pseudo-answer is more accurate and free of irrelevant content. Consequently, all three passages in the second round of utility judgments have utility in answering the question. For the second question in Table 9, the first pseudo-answer is correct, but two passages without utility are judged as “utility”. The second pseudo-answer, with slight rewording, results in all selected passages being correct.
9 Conclusion
In this paper, we propose an Iterative utiliTy judgmEnt fraMework (ITEM) to enhance the utility judgment and QA performance of LLMs by interactions between the steps, inspired by Schutz’s philosophical discussion of relevance. This is a unified framework of iterative RAG with an emphasis on utility. Our framework achieves state-of-the-art performance in zero-shot scenarios, outperforming previous methods in utility judgments, ranking of topical relevance, and answer generation tasks, indicating that the cognitive process of LLMs on a specific topic can also be improved by a similar process. Our experiments also highlight the significance of dynamic interaction in achieving high performance and stability. Future directions include developing better fine-tuning strategies for utility judgments and creating end-to-end solutions for RAG.
Limitations
There are two primary limitations that should be acknowledged: (i) Our methods are applied in zero-shot scenarios without any training. The zero-shot approach itself does not enhance the LLMs’s inherent capability in utility judgments but rather employs strategies to improve performance on utility judgment tasks. Future research should explore designing more effective training methods, e.g., utilizing our iterative framework with self-evolution techniques Singh et al. (2023), to genuinely enhance the LLMs’s ability in utility judgments through training. (ii) The number of candidate passages in the search scenario is much larger than 20. The number of search results we assumed is too small. We need to continue to study utility judgments in large-scale scenarios in the future.
10 Ethics Statement
Our research does not rely on personally identifiable information. All datasets and models used in our paper are publicly available and have been widely adopted by researchers. We firmly believe in the principles of open research and the scientific value of reproducibility. To this end, we have made all data, and code associated with our paper publicly available on GitHub. This transparency not only facilitates the verification of our findings by the community but also encourages the application of our methods in other contexts.
Appendix A Experiment Details
A.1 Effect of Iteration Numbers
The precision, recall, and F1 performance of different LLMs on different datasets with different iteration numbers is shown in Table 6, Table 7, Table 8, and Table 9.
| Method | TREC | WebAP | ||||||||||
| listwise | pointwise | listwise | pointwise | |||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Vanilla | 36.82 | 60.13 | 45.67 | 29.92 | 91.61 | 45.11 | 13.07 | 50.83 | 20.79 | 13.30 | 86.29 | 23.05 |
| UJ-ExpA | 48.51 | 61.15 | 54.10 | 28.12 | 96.27 | 43.53 | 18.83 | 54.16 | 27.94 | 14.65 | 91.82 | 25.27 |
| UJ-ImpA | 40.16 | 60.53 | 48.29 | 33.95 | 83.76 | 48.31 | 16.46 | 52.45 | 25.06 | 17.56 | 73.55 | 28.35 |
| 5-sampling | 46.64 | 59.56 | 52.31 | - | - | - | 20.61 | 56.22 | 30.16 | - | - | - |
| ITEM-As ExpA (=1) | 48.07 | 61.04 | 53.78 | 34.21 | 89.11 | 49.44 | 20.57 | 53.81 | 29.76 | 17.86 | 78.41 | 29.10 |
| ITEM-As ExpA (=2) | 50.58 | 61.86 | 55.65 | 35.87 | 88.73 | 51.09 | 21.11 | 50.85 | 29.83 | 18.27 | 82.00 | 29.88 |
| ITEM-As ExpA (=3) | 50.61 | 59.88 | 54.86 | 36.23 | 90.46 | 51.74 | 23.57 | 48.14 | 31.65 | 18.73 | 81.96 | 30.50 |
| ITEM-As ExpA (=4) | 50.01 | 61.15 | 55.02 | 36.41 | 90.36 | 51.90 | 21.44 | 44.62 | 28.96 | 19.19 | 80.59 | 31.00 |
| ITEM-As ExpA (=5) | 50.61 | 59.88 | 54.86 | 36.14 | 90.46 | 51.65 | 24.07 | 47.09 | 31.86 | 19.17 | 78.94 | 30.85 |
| ITEM-As ImpA (=1) | 39.97 | 64.62 | 49.39 | 30.98 | 89.38 | 46.01 | 16.88 | 57.13 | 26.06 | 17.10 | 81.65 | 28.28 |
| ITEM-As ImpA (=2) | 43.14 | 61.52 | 50.72 | 30.90 | 87.00 | 45.60 | 19.41 | 54.82 | 28.67 | 18.88 | 78.06 | 30.40 |
| ITEM-As ImpA (=3) | 44.43 | 62.82 | 52.05 | 31.68 | 87.99 | 46.59 | 19.21 | 54.20 | 28.36 | 18.69 | 77.77 | 30.13 |
| ITEM-As ImpA (=4) | 44.72 | 61.29 | 51.71 | 31.66 | 87.40 | 46.49 | 17.44 | 47.11 | 25.46 | 18.95 | 78.06 | 30.50 |
| ITEM-As ImpA (=5) | 44.63 | 60.98 | 51.54 | 31.80 | 89.32 | 46.91 | 18.98 | 48.88 | 27.35 | 19.05 | 76.69 | 30.52 |
| ITEM-ARs (=1) | 43.65 | 65.34 | 52.34 | - | - | - | 25.04 | 60.99 | 35.50 | - | - | - |
| ITEM-ARs (=2) | 45.10 | 65.46 | 53.40 | - | - | - | 24.42 | 51.97 | 33.23 | - | - | - |
| ITEM-ARs (=3) | 49.07 | 65.96 | 56.27 | - | - | - | 27.70 | 55.95 | 37.06 | - | - | - |
| ITEM-ARs (=4) | 50.96 | 62.32 | 56.07 | - | - | - | 23.77 | 53.40 | 32.90 | - | - | - |
| ITEM-ARs (=5) | 53.01 | 63.60 | 57.82 | - | - | - | 25.85 | 47.56 | 33.50 | - | - | - |
| Method | TREC | WebAP | ||||||||||
| listwise | pointwise | listwise | pointwise | |||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Vanilla | 34.67 | 85.80 | 49.39 | 31.42 | 98.47 | 47.64 | 12.69 | 77.15 | 21.79 | 14.65 | 87.36 | 25.09 |
| UJ-ExpA | 39.21 | 80.98 | 52.83 | 38.27 | 90.15 | 53.73 | 16.32 | 77.92 | 26.99 | 18.04 | 77.15 | 29.25 |
| UJ-ImpA | 33.92 | 83.36 | 48.22 | 38.68 | 71.47 | 50.20 | 15.57 | 82.79 | 26.22 | 17.22 | 47.61 | 25.29 |
| 5-sampling | 39.04 | 80.98 | 52.68 | - | - | - | 17.52 | 83.49 | 28.97 | - | - | - |
| ITEM-As ExpA (=1) | 39.68 | 82.88 | 53.66 | 37.58 | 84.84 | 52.09 | 17.54 | 63.67 | 27.50 | 19.65 | 74.31 | 31.08 |
| ITEM-As ExpA (=2) | 42.35 | 84.77 | 56.48 | 38.25 | 84.58 | 52.68 | 17.39 | 60.25 | 26.99 | 20.23 | 73.01 | 31.68 |
| ITEM-As ExpA (=3) | 42.00 | 84.15 | 56.03 | 37.84 | 85.50 | 52.46 | 19.12 | 62.87 | 29.32 | 20.91 | 74.63 | 32.67 |
| ITEM-As ExpA (=4) | 41.85 | 84.41 | 55.96 | 38.12 | 85.16 | 52.67 | 17.53 | 61.85 | 27.31 | 20.44 | 73.83 | 32.02 |
| ITEM-As ExpA (=5) | 42.36 | 84.15 | 56.35 | 37.35 | 84.69 | 51.84 | 18.94 | 62.87 | 29.12 | 20.88 | 75.45 | 32.71 |
| ITEM-As ImpA (=1) | 39.63 | 83.42 | 53.73 | 39.70 | 82.87 | 53.68 | 15.48 | 73.66 | 25.59 | 20.04 | 64.06 | 30.53 |
| ITEM-As ImpA (=2) | 38.75 | 85.63 | 53.35 | 38.15 | 82.36 | 52.14 | 15.50 | 76.47 | 25.77 | 18.54 | 62.69 | 28.62 |
| ITEM-As ImpA (=3) | 40.84 | 84.86 | 55.14 | 40.58 | 79.64 | 53.76 | 15.99 | 70.99 | 26.10 | 19.54 | 61.32 | 29.64 |
| ITEM-As ImpA (=4) | 38.88 | 82.74 | 52.90 | 39.34 | 81.74 | 53.12 | 15.03 | 74.41 | 25.01 | 19.72 | 59.95 | 29.68 |
| ITEM-As ImpA (=5) | 41.26 | 84.61 | 55.47 | 40.92 | 82.14 | 54.63 | 15.49 | 68.93 | 25.29 | 19.84 | 57.21 | 29.46 |
| ITEM-ARs (=1) | 34.53 | 84.17 | 48.97 | - | - | - | 20.05 | 72.88 | 31.44 | - | - | - |
| ITEM-ARs (=2) | 36.27 | 83.19 | 50.51 | - | - | - | 15.92 | 79.01 | 26.50 | - | - | - |
| ITEM-ARs (=3) | 38.04 | 82.68 | 52.10 | - | - | - | 17.93 | 76.87 | 29.08 | - | - | - |
| ITEM-ARs (=4) | 37.28 | 83.70 | 51.58 | - | - | - | 16.60 | 78.81 | 27.42 | - | - | - |
| ITEM-ARs (=5) | 40.25 | 81.37 | 53.86 | - | - | - | 17.04 | 74.83 | 27.75 | - | - | - |
| Method | TREC | WebAP | ||||||||||
| listwise | pointwise | listwise | pointwise | |||||||||
| P | R | F1 | P | R | F1 | P | R | F1 | P | R | F1 | |
| Vanilla | 42.13 | 79.98 | 55.19 | 33.86 | 94.40 | 49.84 | 17.13 | 83.45 | 28.43 | 15.80 | 89.42 | 26.85 |
| UJ-ExpA | 45.74 | 77.36 | 57.49 | 32.06 | 96.19 | 48.09 | 19.51 | 69.86 | 30.50 | 16.23 | 88.74 | 27.44 |
| UJ-ImpA | 44.19 | 77.11 | 56.18 | 33.45 | 90.36 | 48.83 | 18.37 | 80.14 | 29.89 | 15.58 | 84.51 | 26.32 |
| 5-sampling | 50.78 | 74.77 | 60.49 | - | - | - | 20.70 | 65.83 | 31.49 | - | - | - |
| ITEM-As ExpA (=1) | 55.55 | 71.48 | 62.52 | 37.83 | 91.94 | 53.61 | 26.74 | 59.45 | 36.89 | 19.73 | 84.95 | 32.02 |
| ITEM-As ExpA (=2) | 57.95 | 70.40 | 63.57 | 40.74 | 93.04 | 56.67 | 29.43 | 60.58 | 39.62 | 19.62 | 78.62 | 31.40 |
| ITEM-As ExpA (=3) | 58.36 | 68.88 | 63.18 | 40.00 | 91.88 | 55.74 | 29.30 | 60.91 | 39.57 | 19.80 | 76.20 | 31.43 |
| ITEM-As ExpA (=4) | 58.48 | 70.67 | 64.00 | 40.25 | 93.38 | 56.25 | 29.11 | 61.03 | 39.42 | 20.48 | 79.63 | 32.58 |
| ITEM-As ExpA (=5) | 58.34 | 69.69 | 63.51 | 39.29 | 92.16 | 55.09 | 29.76 | 60.68 | 39.93 | 20.58 | 80.42 | 32.77 |
| ITEM-As ImpA (=1) | 54.36 | 65.08 | 59.24 | 40.89 | 82.20 | 54.61 | 24.79 | 64.37 | 35.80 | 18.78 | 67.00 | 29.34 |
| ITEM-As ImpA (=2) | 55.88 | 63.11 | 59.27 | 43.32 | 83.13 | 56.96 | 27.68 | 62.03 | 38.28 | 20.70 | 70.54 | 32.00 |
| ITEM-As ImpA (=3) | 57.33 | 64.17 | 60.56 | 41.66 | 80.48 | 54.90 | 30.01 | 63.60 | 40.78 | 21.51 | 66.77 | 32.54 |
| ITEM-As ImpA (=4) | 55.98 | 62.24 | 58.95 | 42.34 | 80.65 | 55.53 | 28.43 | 60.11 | 38.60 | 20.60 | 65.63 | 31.36 |
| ITEM-As ImpA (=5) | 56.63 | 62.19 | 59.28 | 41.49 | 83.57 | 55.45 | 29.05 | 60.66 | 39.29 | 21.51 | 68.03 | 32.68 |
| ITEM-ARs (=1) | 51.94 | 76.90 | 62.00 | - | - | - | 25.32 | 65.84 | 36.58 | - | - | - |
| ITEM-ARs (=2) | 53.77 | 76.19 | 63.05 | - | - | - | 25.55 | 59.26 | 35.70 | - | - | - |
| ITEM-ARs (=3) | 52.41 | 74.04 | 61.37 | - | - | - | 27.61 | 63.96 | 38.58 | - | - | - |
| ITEM-ARs (=4) | 52.75 | 73.78 | 61.52 | - | - | - | 28.84 | 61.85 | 39.34 | - | - | - |
| ITEM-ARs (=5) | 52.77 | 76.28 | 62.39 | - | - | - | 28.76 | 62.54 | 39.40 | - | - | - |
| References of RAG | Mistral | Llama 3 | ChatGPT | |||
| EM | F1 | EM | F1 | EM | F1 | |
| Golden Evidence | 46.09 | 62.59 | 64.45 | 76.64 | 66.40 | 76.86 |
| RocketQAv2 | 31.58 | 47.69 | 50.96 | 62.01 | 46.54 | 57.00 |
| Vanilla | 31.16 | 47.43 | 49.09 | 60.56 | 48.52 | 58.64 |
| UJ-ExpA | 32.76 | 48.46 | 49.63 | 61.10 | 47.72 | 58.01 |
| UJ-ImpA | 30.67 | 46.83 | 48.88 | 60.26 | 49.01 | 59.30 |
| 5-sampling | 33.24 | 48.84 | 48.72 | 60.71 | 48.90 | 58.97 |
| ITEM-As ExpA (=1) | 32.98 | 49.00 | 50.16 | 61.88 | 49.38 | 59.78 |
| ITEM-As ExpA (=2) | 34.31 | 50.08 | 50.48 | 62.32 | 49.22 | 59.99 |
| ITEM-As ExpA (=3) | 33.73 | 49.63 | 50.27 | 62.09 | 49.69 | 60.18 |
| ITEM-As ExpA (=4) | 34.21 | 50.07 | 50.43 | 62.20 | - | - |
| ITEM-As ExpA (=5) | 33.78 | 49.63 | 50.27 | 62.07 | - | - |
| ITEM-As ImpA (=1) | 32.17 | 48.51 | 50.37 | 61.89 | 48.75 | 58.99 |
| ITEM-As ImpA (=2) | 32.49 | 48.67 | 49.63 | 61.16 | 49.11 | 59.14 |
| ITEM-As ImpA (=3) | 32.39 | 48.47 | 49.68 | 61.48 | 48.69 | 58.94 |
| ITEM-As ImpA (=4) | 32.71 | 48.84 | 49.41 | 61.03 | - | - |
| ITEM-As ImpA (=5) | 32.33 | 48.44 | 49.73 | 61.42 | - | - |
| ITEM-ARs (=1) | 33.30 | 49.26 | 50.27 | 61.69 | 49.52 | 59.64 |
| ITEM-ARs (=2) | 33.57 | 49.16 | 50.70 | 61.92 | 49.01 | 59.75 |
| ITEM-ARs (=3) | 33.40 | 49.27 | 49.36 | 60.97 | 49.06 | 59.67 |
| ITEM-ARs (=4) | 33.46 | 49.24 | 49.84 | 61.54 | - | - |
| ITEM-ARs (=5) | 33.89 | 49.58 | 49.20 | 60.84 | - | - |
A.2 Quality of Utility Judgments
The relevance labels of TREC DL are of a four-point scale, and we consider the highest level as having utility. To see the utility judgment performance when we consider lower grades to have utility, we measure the precision of utility judgments of Mistral on TREC DL when passages of different grades are treated as positive in Table 10.
| label1 | label2 | label3 | |
| =1 | 82.08 | 68.34 | 48.07 |
| =2 | 83.86 | 69.53 | 50.58 |
| =3 | 84.23 | 71.06 | 50.61 |
| =4 | 84.63 | 70.18 | 50.01 |
| =5 | 84.52 | 70.69 | 50.61 |
We can see that almost 70% of the results of positive utility judgments are highly relevant to the question.
A.3 values in ITEM-Ar
Different ranking performance of values in ITEM-Ar is shown in Table 11. Considering the performance of utility ranking and utility judgments, we set =5.
| , | Ranking | Utility judgments | ||||||
| N@1 | N@3 | N@5 | N@10 | N@20 | P | R | F1 | |
| =1, =1 | 72.76 | 71.27 | 70.57 | 72.69 | 84.08 | 53.66 | 24.09 | 33.25 |
| =1, =2 | 76.02 | 71.54 | 71.38 | 73.66 | 84.78 | 58.54 | 28.73 | 38.54 |
| =1, =3 | 77.24 | 72.83 | 71.83 | 73.87 | 85.20 | 59.76 | 28.84 | 38.90 |
| =1, =4 | 77.24 | 73.04 | 71.91 | 73.90 | 85.25 | 59.76 | 28.84 | 38.90 |
| =1, =5 | 76.02 | 72.11 | 71.42 | 73.45 | 84.98 | 58.54 | 28.71 | 38.53 |
| =5, =1 | 72.76 | 71.27 | 70.57 | 72.69 | 84.08 | 33.17 | 57.31 | 42.02 |
| =5, =2 | 78.46 | 73.74 | 72.86 | 75.48 | 86.09 | 32.93 | 58.37 | 42.10 |
| =5, =3 | 79.27 | 75.00 | 74.27 | 75.78 | 86.80 | 34.15 | 62.57 | 44.18 |
| =5, =4 | 79.67 | 75.92 | 75.35 | 76.83 | 87.23 | 35.12 | 61.40 | 44.68 |
| =5, =5 | 79.67 | 75.32 | 74.61 | 76.20 | 86.82 | 34.63 | 61.25 | 44.25 |
| =10, =1 | 72.76 | 71.27 | 70.57 | 72.69 | 84.08 | 22.56 | 68.03 | 33.88 |
| =10, =2 | 78.05 | 72.64 | 72.90 | 75.48 | 85.74 | 23.66 | 75.47 | 36.02 |
| =10, =3 | 80.89 | 76.58 | 74.54 | 76.30 | 86.94 | 23.78 | 75.65 | 36.19 |
| =10, =4 | 78.05 | 74.70 | 72.85 | 75.12 | 85.72 | 24.51 | 74.17 | 36.85 |
| =10, =5 | 79.67 | 75.60 | 74.84 | 76.54 | 86.88 | 23.66 | 74.42 | 35.90 |
A.4 ITEM-ARr
We evaluate two ranking performances of ITEM-ARr during the same loop, with the experimental results shown in Table 12. We find that under the ITEM-ARr framework, relevance ranking and utility ranking are both improved, and utility ranking performance is generally better than relevance ranking. However, as seen in Table 3 and Table 12, performing ranking twice in the same iteration may not yield better ranking results than performing utility ranking once in the iteration.
| NDCG@5 | NDCG@10 | NDCG@20 | Utility-F1 | |
| 1 | 71.29 / 72.77 | 72.90 / 74.96 | 84.56 / 85.75 | 43.13 |
| 2 | 72.54 / 70.99 | 74.81 / 73.76 | 85.77 / 85.28 | 40.21 |
| 3 | 72.07 / 74.14 | 74.14 / 76.63 | 85.53 / 86.57 | 45.67 |
| 4 | 71.02 / 71.06 | 74.30 / 74.03 | 85.09 / 85.16 | 43.82 |
| 5 | 72.26 / 70.12 | 75.83 / 72.59 | 85.88 / 84.77 | 44.10 |
Appendix B Case Study
We show two cases on the TREC dataset in Table 9.
Appendix C Datasets and Evaluation
| Dataset | #Psg | #PsgLen | #Q | #Rels/Q |
| TREC | 8.8M | 93 | 82 | 212.8 |
| WebAP | 379k | 45 | 73 | 76.4 |
| NQ | 21M | 100 | 1868 | 1.0 |
| GTI-NQ | 10 | 100 | 1863 | 1.0 |
Detailed statistics of the experimental datasets are shown in Table 13. We use the official tool for evaluation of NDCG.
Appendix D Answer Passage Retrieval
Non-factoid questions are usually expected longer answers, such as sentence-level or passages-level Keikha et al. (2014a); Yang et al. (2016); Keikha et al. (2014b). Yang et al. (2016) developed an annotated dataset for answer passage retrieval called WebAP, which has an average of 76.4 qrels per query. Cohen et al. (2018) and Hashemi et al. (2020) introduced the WikiPassageQA dataset and ANTIQUE dataset for answer passage retrieval, respectively. Compared to the WebAP dataset, WikiPassageQA and ANTIQUE have incomplete annotations, with an average of 1.7 qrels and 32.9 qrels per query Hashemi et al. (2019, 2020). Bi et al. (2019) created the PsgRobust dataset for answer passage retrieval, which is built on the TREC Robust collection Voorhees et al. (2003) following a similar approach to WebAP but without manual annotation.
Appendix E -sampling
The output of -sampling each time contains explicit answers and utility judgments. If the question length is , the total length of the input passages is , and the average length of a single passage is , then the k-sampling input cost is . If the average length of the output explicit answer is , and the average length of the output utility judgments is , then the k-sampling output cost is . Taking ITEM-As as an example, with a maximum of three iterations, the maximum input cost for utility judgments is . For answer generation, the longest input is and the shortest is . Therefore, the maximum input cost for ITEM-As is and the minimum is . The maximum output cost is . Therefore, in order to ensure fairness in the calculation of large language model parameters, we choose =5.
Appendix F Retrievers
We use two representative retrievers to gather candidate passages in for utility judgments. Following with previous works Zhang et al. (2024); Sun et al. (2023), we employ RocketQAv2 Ren et al. (2021) and BM25 Robertson et al. (2009) for the NQ dataset and retrieval datasets(i.e., TREC DL and WebAP datasets), respectively. Based on the retrieval results to build the we have two settings: (i) For TREC DL and WebAP datasets, we select the top-20 BM25 retrieval results. If these do not include passages with utility, we replaced the last one with a utility-annotated passage. (ii) For the NQ dataset, we use the top-10 dense retrieval results to form the candidate list , following the GTU setting of Zhang et al. (2024).
Appendix G Inference Efficiency
Table 14 shows more analysis of Mistral on the TREC DL dataset using the listwise input form. The temperature of the LLMs is set to 0 during the generation process, and we used a single run.
| Methods | #IT/Q | #OT/Q | #RT(s)/Q |
| Vanilla | 2507 | 23 | 0.49 |
| UJ-ExpA | 2529 | 59 | 0.94 |
| UJ-ImpA | 2533 | 41 | 0.71 |
| 5-sampling | 12647 | 296 | 4.72 |
| ITEM-As(m=1) | 4628 | 47 | 0.96 |
| ITEM-As(m=3) | 10603 | 154 | 2.13 |
| ITEM-ARs(m=1) | 7107 | 211 | 3.15 |
| ITEM-ARs(m=3) | 18224 | 624 | 8.61 |
Due to the iterations, the average input token length of our methods is relatively large. The cost of ITEM-As is about 0.5 times that of the 5-sampling, and ITEM-ARs is about 1.5 times the cost of 5-sampling. Our framework provides a way of automatically obtaining high-quality labeled data for each task in RAG. These annotations can be used to train regular RAG models. The cost should be worth it (much better than -sampling).
Appendix H Instruction Details
H.1 Instruction of Listwise and Pointwise Approaches
For the prompts of the NQ dataset using ChatGPT, we follow the setting of Zhang et al. (2024), otherwise, we use the following prompts. Following Sun et al. (2023), we input N passages using the form of multiple rounds of dialogue in the listwise approach. The prompts we used in our experiments are shown in Figure 10 and Figure 11.
H.2 Instruction of the Ranking Approach
For RankGPT, we directly use the instruction of Sun et al. (2023) for relevance ranking, as shown in Figure 14. For the relevance ranking in our ITEM, the instructions are shown in Figure 12 and Figure 13.
H.3 Instruction of Answer Generation
Li et al. (2023) utilize LLM to generate the missing information in the provided documents for the current question and then re-retrieve it as relevant feedback. Therefore, we have also designed two kinds of pseudo answers for utility judgments, i.e., (i) the explicit answer, which produces an answer based on the given information, and (ii) the implicit answer, which does not answer the question directly but gives the information necessary to answer the question. The two instructions are shown in Figure 15 and Figure 16.